### v2

# On variants of Chebyshev’s conjecture

Abstract.
We show that the sign constancy for the values of certain weighted summatory functions of the von Mangoldt function implies the Riemann hypothesis or the generalized Riemann hypothesis for Dirichlet $L$-functions. While such sign constancy is challenging to establish individually, we prove that the summatory functions under study have constant signs on average.
Key words and phrases:
Chebyshev’s bias, Riemann Hypothesis, Generalized Riemann Hypothesis, screw functions
Mathematics Subject Classification:
11M26 11N05
1. Introduction
1.1. Results on the Riemann zeta function
Let $\pi(x)$ be the number of prime numbers $\leq x$. According to the prime number theorem, the leading term in the asymptotic behavior of $\pi(x)$ as $x\to\infty$ is the logarithmic integral ${\rm li}(x)=\int_{0}^{x}dt/\log t$. The difference $\pi(x)-{\rm li}(x)$ had long been conjectured to be negative for all $x$. This conjecture was disproven in 1914 by Littlewood, who showed that it changes sign infinitely often [8]. However, Rubinstein and Sarnak [19] proved that the logarithmic density of $x$ for which $\pi(x)-{\rm li}(x)$ is negative is close to one assuming the Riemann Hypothesis (RH) holds – that is, all nontrivial zeros of $\zeta(s)$ have real part $1/2$ – and the set of positive imaginary parts of the nontrivial zeros of the Riemann zeta function $\zeta(s)$ is linearly independent over $\mathbb{Q}$. Furthermore, as mentioned in Pintz [16] and as proved by Johnston [10], the RH is equivalent to the negativity
$\int_{2}^{x}(\pi(t)-{\rm li}(t))\,dt<0\quad(x>2).$ (1.1)
Johnston also shows that the RH is equivalent to the negativity
$\int_{2}^{x}(\vartheta(t)-t)\,dt<0\quad(x>2),$ (1.2)
where $\vartheta(x)=\sum_{p\leq x}\log p$, the first Chebyshev function. Here and in what follows, the letter $p$ will always denote a prime number. In contrast, no such bias is observed in the difference between the second Chebyshev function
$\psi(x):=\sideset{}{{}^{\prime}}{\sum}_{n\leq x}\Lambda(n)$
and its leading term $x$ as $x\to\infty$, where $\Lambda(n)$ is the von Mangoldt function defined by $\Lambda(n)=\log p$ if $n=p^{k}$ with $k\in\mathbb{Z}_{>0}$ and $\Lambda(n)=0$ otherwise, and $\sideset{}{{}^{\prime}}{\sum}_{n\leq x}a_{n}$ means $\sum_{n\leq x}a_{n}-(1/2)a_{x}$ when $x$ is a prime power and $\sum_{n\leq x}a_{n}$ otherwise.
On the other hand, a negative bias is observed in the difference between the weighted Chebyshev function
$\psi_{1/2}(x):=\sideset{}{{}^{\prime}}{\sum}_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}$ (1.3)
and its leading term $2\sqrt{x}$ as $x\to\infty$. Regarding this observation, using the explicit formula
$\psi_{1/2}(x)-2\sqrt{x}=-\sum_{\rho}\frac{x^{\rho-1/2}}{\rho-1/2}-\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)+o(1)$ (1.4)
in (2.6) below, together with Littlewood’s result [8, Theorem 5.6]
$\sum_{\rho}\frac{x^{\rho}}{\rho}=\Omega_{\pm}(\sqrt{x}\log\log\log x),$
we can show that $\psi_{1/2}(x)-2\sqrt{x}$ changes sign infinitely often as well as $\pi(x)-{\rm li}(x)$, where the sum $\sum_{\rho}$ runs over all nontrivial zeros of $\zeta(s)$ counted with multiplicity, and is defined as $\sum_{\rho}=\lim_{T\to\infty}\sum_{|\Im(\rho)|\leq T}$. The Landau symbol $f(x)=o(g(x))$ means that $|f(x)|/g(x)$ tends to zero as $x\to\infty$. Furthermore, $f(x)=\Omega_{+}(g(x))$ (resp. $f(x)=\Omega_{-}(g(x))$) means that $\limsup_{x\to\infty}f(x)/g(x)>0$ (resp. $\liminf_{x\to\infty}f(x)/g(x)<0$), and $f(x)=\Omega_{\pm}(g(x))$ means that both $f(x)=\Omega_{+}(g(x))$ and $f(x)=\Omega_{-}(g(x))$ hold. By the second functional equation in (2.4), if $\rho$ is a nontrivial zero of $\zeta(s)$, then so is $\bar{\rho}$. Therefore, both $\sum_{\rho}x^{\rho-1/2}/\rho$ and $\sum_{\rho}x^{\rho-1/2}/(\rho-1/2)$ are real valued.
However, assuming the RH, $\psi_{1/2}(x)-2\sqrt{x}$ is negative on average, because
$\int_{0}^{x}(\psi_{1/2}(y)-2\sqrt{y}\,)\frac{dy}{y}=-\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)\log x-\sum_{\rho}\frac{x^{\rho-1/2}}{(\rho-1/2)^{2}}+o(1)$
holds, $(\zeta^{\prime}/\zeta)(1/2)=2.68609\dots$ (see (3.1) and (3.2) below), and the sum on the right-hand side is bounded under the RH by $\sum_{\rho}|\rho|^{-2}<\infty$ (see §2.2). This bias in the difference $\psi_{1/2}(x)-2\sqrt{x}$ is also a sufficient condition for the RH as in the case of $\pi(x)-{\rm li}(x)$ and $\vartheta(x)-x$:
Theorem 1.
The following three statements are equivalent:
  1. (1)
    The RH holds.
  2. (2)
    There exists an $x_{0}\geq 2$ such that
    $\int_{0}^{x}$ $(\psi_{1/2}(y)-2\sqrt{y}\,)\frac{dy}{y}$ (1.5)
    $=\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}-4\sqrt{x}~\leq~0$
    holds for all $x\geq x_{0}$.
  3. (3)
    It holds that
    $\lim_{x\to\infty}\left[\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)-\frac{4\sqrt{x}}{\log x}\right]=-\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right).$ (1.6)
The prime on the sum is unnecessary on the left-hand side of (1.5), since $\log(x/n)$ vanishes when $n=x$. Taking the average
$\frac{1}{x}\int_{1}^{x}\frac{1}{2}\sqrt{\frac{x}{u}}\log\sqrt{\frac{x}{u}}\,du=1-\frac{1}{2\sqrt{x}}(\log x+2)$
into account, the inequality in (1.5) can be rewritten as an inequality for the sum of $\Lambda(n)$ with a weight that is approximately one on average:
$\sum_{n\leq x}\Lambda(n)\cdot\frac{1}{2}\sqrt{\frac{x}{n}}\log\sqrt{\frac{x}{n}}-x~\leq~0.$
This representation is sometimes used for its apparent simplicity as in Theorem 4 below.
Considering [19], we are interested in the logarithmic density of $x$ for which $\psi_{1/2}(x)-2\sqrt{x}\leq 0$ holds, but we do not study that topic in this paper.
The equivalence between the first and second statements in Theorem 1 is quite similar to that between the RH and either (1.1) or (1.2). However, an analog of (1.6) for $\pi(x)-\mathrm{li}(x)$ or $\vartheta(x)-x$ does not appear to have been previously noted in the literature. Moreover, it should be mentioned that Theorem 1 was not discovered as an analog of (1.1); rather, it arose in connection with the screw functions discussed below.
Theorem 1 can be interpreted in terms of the screw function $g_{\zeta}(t)$ for $\zeta(s)$ studied in [21]. It is a function on $\mathbb{R}$ defined by
$g_{\zeta}(t):=g_{0}(t)+g_{\infty}(t)$
with
$g_{0}(t)$ $=\sum_{n\leq\exp(|t|)}\frac{\Lambda(n)}{\sqrt{n}}(|t|-\log n)-4(e^{t/2}+e^{-t/2}-2),$ (1.7)
$g_{\infty}(t)$ $=-\frac{|t|}{2}\left(\frac{\Gamma^{\prime}}{\Gamma}\left(\frac{1}{4}\right)-\log\pi\right)-\frac{1}{4}\Bigl{(}\Phi(1,2,1/4)-e^{-|t|/2}\Phi(e^{-2|t|},2,1/4)\Bigr{)},$
where $\Phi(z,s,a)=\sum_{n=0}^{\infty}z^{n}(n+a)^{-s}$ is the Hurwitz–Lerch zeta function for $|z|<1$ and $a\not=0,-1,-2,\cdots$. It is known that $g_{\zeta}(t)\geq 0$ for all $t\in\mathbb{R}$ if and only if the RH holds [21, Theorem 1.7]. The nonnegativity comes from the fact that $g_{\zeta}(t)$ is a screw function in the sense of [12, p. 189] under the RH. From the proof of the integral formula
$\int_{0}^{\infty}g_{\zeta}(t)e^{izt}\,dt=\frac{1}{z^{2}}\frac{\xi^{\prime}}{\xi}\left(\frac{1}{2}-iz\right)$
in [21, Section 2.1], the function $g_{0}(t)$ corresponds to the non-archimedean factor $s(s-1)\zeta(s)$ and $g_{\infty}(t)$ corresponds to the archimedean factor $\pi^{-s/2}\Gamma(s/2)$ of the Riemann xi-function
$\xi(s):=\frac{1}{2}\,s(s-1)\pi^{-s/2}\Gamma(s/2)\zeta(s).$ (1.8)
According to the Kreĭn–Langer criterion [12, Satz 5.9], $g_{0}(t)$ cannot be a screw function even assuming the RH, because the logarithmic derivative of $s(s-1)\zeta(s)$ can never belong to the Nevanlinna class. However, if we note the relation
$g_{0}(\log x)=\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}-4\left(\sqrt{x}+\frac{1}{\sqrt{x}}-2\right),$ (1.9)
then we find that the RH is equivalent to $g_{0}(t)$ having a constant sign.
Corollary 1.
Let $g_{0}(t)$ be the non-archimedean part of $g_{\zeta}(t)$ defined in (1.7). Then, the RH holds if and only if there exists $t_{0}>0$ such that $g_{0}(t)\geq 0$ holds for all $t\geq t_{0}$.
Noting (1.4), if we replace $4\sqrt{x}$ in (1.5) with $2\sum_{n\leq x}\Lambda(n)/\sqrt{n}$ and express $2$ as $\log\exp(2)$, we obtain the following as an analog of Theorem 1.
Theorem 2.
Suppose that there exists an $x_{0}\geq 2$ such that
$\sum_{n\leq xe^{2}}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}~\leq~0$ (1.10)
holds for all $x\geq x_{0}$. Then, the RH holds. Furthermore, it holds that
$\lim_{x\to\infty}\sum_{n\leq xe^{2}}\frac{\Lambda(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)=-\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)$ (1.11)
if and only if the RH holds and the estimate
$\sum_{\rho}\frac{x^{\rho}}{\rho}=o(\sqrt{x}\log x)$ (1.12)
is valid as $x\to\infty$.
The right-hand side of (1.11) is negative (see (3.2) below). Therefore, (1.11) implies that there exists $x_{0}\geq 2$ such that (1.10) holds for all $x\geq x_{0}$,
For the sum on the left-hand side of (1.12), the best known estimate under the RH is
$\sum_{\rho}\frac{x^{\rho}}{\rho}\ll\sqrt{x}(\log x)^{2}\quad\text{as $x\to\infty$}$
[9, pp. 82–85], but it is not enough to prove (1.10) for large $x$, where the Vinogradov symbol $f(x)\ll g(x)$ mean that there exists a positive constant $M$ such that $|f(x)|\leq Mg(x)$ holds for a prescribed range of $x$. We use the the Landau symbol $f(x)=O(g(x))$ in the same meaning. In the above sense, (1.10) and (1.11) may provide a condition stronger than the RH. For comparison, we note that Montgomery’s conjecture
$\limsup_{x\to\infty}\frac{\sum_{\rho}x^{\rho}/\rho}{\sqrt{x}(\log\log\log x)^{2}}=\frac{1}{2\pi}\quad\text{and}\quad\liminf_{x\to\infty}\frac{\sum_{\rho}x^{\rho}/\rho}{\sqrt{x}(\log\log\log x)^{2}}=-\frac{1}{2\pi}$
[14, p.16] implies the RH and the estimate (1.12). The equivalence of (1.11) and (1.12) can be viewed as analogous to Akatsuka [2, Theorem 1], which in turn parallels Conrad [5, Theorem 1.3] for $\zeta(s)$. Taking the logarithm of [2, Corollary 4.4] and using [2, Lemma 2.1], Akatsuka’s result asserts that
$\lim_{x\to\infty}\left[\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}\log n}-\lim_{\varepsilon\to 0+}\left(\int_{0}^{1-\varepsilon}+\int_{1+\varepsilon}^{\sqrt{x}}\right)\frac{du}{\log u}\right]=\log\left(-\zeta\left(\frac{1}{2}\right)\right)$
is equivalent to the validity of the RH and (1.12), where $\zeta(1/2)=1.46035\cdots$. Recently, Akatsuka showed that the equation obtained by replacing the upper limit $\sqrt{x}$ of the integral on the left-hand side with $\sqrt{\vartheta(x)}$ is equivalent to the RH ([3, Theorem 3]).
A sequence $\{a_{n}\}$ is called Riesz summable of order $k$ to $a$ if
$\lim_{x\to\infty}\sum_{n\leq x}a_{n}\left(1-\frac{\log n}{\log x}\right)=a$
(cf. [15, p.158]). From (3.1) below, $\{\Lambda(n)/\sqrt{n}\}$ cannot be Riesz summable of order one even if the RH is assumed. However, Theorem 2 asserts that if we suppose a condition stronger than the RH, then a slightly modified Riesz sum of order one has a limit, and conversely, the existence of such a limit implies the RH.
If we rewrite (1.10) as
$\sum_{n\leq x}\Lambda(n)\sqrt{\frac{x}{n}}\log\frac{x}{n}\quad\leq~\sum_{x<n\leq xe^{2}}\Lambda(n)\sqrt{\frac{x}{n}}\log\frac{n}{x}$ (1.13)
it could be interpreted as a bias depending on the size of the weighted prime powers.
1.2. Results related to Chebyshev’s conjecture
For any pair of integers $q$ and $a$ with $(q,a)=1$, let $\pi(x;q,a)$ denote the number of primes $p\leq x$ such that $p\equiv a$ mod $q$. The leading term in the asymptotic behavior of $\pi(x,q,a)$ as $x\to\infty$ is ${\rm li}(x)/\varphi(q)$, where $\varphi(q)$ is the Euler totient function. As in the case of the difference $\pi(x)-{\rm li}(x)$, Chebyshev’s conjecture that $\pi(x;4,1)-\pi(x;4,3)$ is always nonpositive was disproved in [8], but it has been shown that the logarithmic density of the set of $x>2$ for which $\pi(x;4,1)-\pi(x;4,3)$ is negative is close to one [19].
Now, considering the similarity between $\pi(x)-{\rm li}(x)$ and $\pi(x;4,1)-\pi(x;4,3)$, we introduce the weighted function
$\psi_{1/2}(x;q,a):=\sideset{}{{}^{\prime}}{\sum}_{{n\leq x}\atop{n\equiv a\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}$
for integers $q$ and $a$ with $(q,a)=1$. Then, a negative bias is observed in the difference
$\psi_{1/2}(x;4,1)-\psi_{1/2}(x;4,3)=\sideset{}{{}^{\prime}}{\sum}_{n\leq x}\frac{\Lambda(n)\chi_{4}(n)}{\sqrt{n}},$
where $\chi_{4}$ is the non-principal Dirichlet character modulo four. Furthermore, we find that $\psi_{1/2}(x;4,1)-\psi_{1/2}(x;4,3)$ is negative on average if and only if all zeros of the Dirichlet $L$-function $L(s,\chi_{4})$ in the critical strip $0\leq\Re(s)\leq 1$ lie on the line $\Re(s)=1/2$. The latter is nothing but the Generalized Riemann Hypothesis (GRH) for $L(s,\chi_{4})$ holds.
Theorem 3.
The following three statements are equivalent:
  1. (1)
    The GRH for $L(s,\chi_{4})$ holds.
  2. (2)
    There exists an $x_{0}\geq 2$ such that
    $\int_{0}^{x}(\psi_{1/2}(y;4,1)$ $-\psi_{1/2}(y;4,3))\frac{dy}{y}$ (1.14)
    $=\sum_{n\leq x}\frac{\Lambda(n)\chi_{4}(n)}{\sqrt{n}}\log\frac{x}{n}~\leq~0$
    holds for all $x\geq x_{0}$.
  3. (3)
    It holds that
    $\lim_{x\to\infty}\sum_{n\leq x}\frac{\Lambda(n)\chi_{4}(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)=-\frac{L^{\prime}}{L}\left(\frac{1}{2},\chi_{4}\right).$ (1.15)
Unlike [5, Theorem 1.3], [2, Theorem 1], and Theorem 2, which include claims stronger than the GRH, it should be noted that (1.15) is equivalent to the GRH for $L(s,\chi_{4})$. On the other hand, similar to Theorem 1, the nonpositivity in (1.14) can be interpreted as a property of the non-archimedean part of the screw function associated with $L(s,\chi_{4})$ studied in [22], but we omit the details, as they may be easily inferred.
Based on his observation of the negative bias in the difference $\pi(x;4,1)-\pi(x;4,3)$, Chebyshev conjectured
$\lim_{x\to\infty}\sum_{p>2}(-1)^{(p-1)/2}\cdot e^{-p/x}=-\infty.$ (1.16)
The sum on the left-hand side of (1.16) can be written as
$\sum_{p\equiv 1\,{\rm mod}\,4}e^{-p/x}\quad-\sum_{p\equiv 3\,{\rm mod}\,4}e^{-p/x}.$
Here $e^{-p/x}\to 1$ as $x\to\infty$ termwise. Therefore, Chebyshev’s conjecture (1.16) suggests that there are “more” primes of the form $p=4n+3$ than those of the form $p=4n+1$. Subsequently, Hardy and Littlewood [8] and Landau [13] proved that (1.16) and
$\lim_{x\to\infty}\sum_{p>2}(-1)^{(p-1)/2}\log p\cdot e^{-p/x}=$-$\infty$ (1.17)
are equivalent to the GRH for $L(s,\chi_{4})$. Knapowski and Turán [11] showed that replacing the weight function $\exp(-p/x)$ in (1.16) and (1.17) with $\exp(-(\log(p/x))^{2})$ is likewise equivalent to the GRH for $L(s,\chi_{4})$. Fujii [7] proved that when $\exp(-p/x)$ is replaced with $\exp(-(p/x)^{\alpha})$ (resp. $2K_{0}(2\sqrt{p/x})$), the result in (1.16) (resp. (1.17)) is equivalent to the GRH for $L(s,\chi_{4})$ if $0<\alpha<4.19$, and he conjectured that this would hold for general $\alpha>0$ as well. Platt and Trudgian [18] extends Fujii’s result for (1.16) to $0<\alpha<20.4$. Aoki and Koyama [1] studies the case where $\exp(-p/x)$ is replaced by $p^{-s}$ ($s\geq 0$) in (1.16). In addition, a few other weight functions were considered in the works of Hardy–Littlewood [8] and Fujii [7]. Based on the above developments and Theorem 3, considering the weight function $\sqrt{x/p}\log(x/p)$ in (1.17) leads to the following result.
Theorem 4.
The GRH for $L(s,\chi_{4})$ holds if and only if
$\lim_{x\to\infty}\sum_{2<p\leq x}(-1)^{(p-1)/2}\log p\cdot\sqrt{\frac{x}{p}}\,\log\frac{x}{p}=-\infty.$ (1.18)
As a result comparable to Theorem 4, we note that the asymptotic formula
$\sum_{2<p\leq x}(-1)^{(p-1)/2}\sqrt{\frac{x}{p}}~\sim~\frac{1}{2}\sqrt{x}\log\log x\quad\text{as $x\to\infty$}$ (1.19)
holds if and only if the GRH for $L(s,\chi_{4})$ holds and (1.12) is satisfied for the nontrivial zeros of $L(s,\chi_{4})$ by combining [1, Theorem 1.1, (1.3)] and [5, Theorem 6.2]. According to the proof of Theorem 4, the GRH for $L(s,\chi_{4})$ is equivalent to the left-hand side of (1.18) being asymptotically equal to $(\sqrt{x}/4)(\log x)^{2}$; thus, unlike (1.19), it does not yield a result stronger than the GRH. However, an analog of Theorem 4 for $\zeta(s)$ yields a conclusion stronger than the RH, similar to Theorem 2.
Theorem 5.
The RH holds assuming
$\lim_{x\to\infty}\sum_{p\leq xe^{2}}\log p\cdot\sqrt{\frac{x}{p}}\,\log\frac{x}{p}=-\infty.$ (1.20)
Conversely, if the RH holds and (1.12) is valid as $x\to\infty$, then (1.20) holds.
1.3. Results on averages over Dirichlet characters
As naturally expected, Theorem 3 generalizes to a result for general Dirichlet $L$-functions, as stated in Theorem 8 below. This generalization is also described by inequalities for oscillating sums, however, these sums can be rewritten as those with a constant sign in both the first and second halves. For the following results, for an imprimitive Dirichlet character $\chi$, we define the GRH to hold for $L(s,\chi)$ as the GRH to hold for $L(s,\chi^{\ast})$, where $\chi^{\ast}$ is a primitive Dirichlet character inducing $\chi$.
Theorem 6.
Let $q$ be a positive integer, and let $\beta=\beta(q)$ be a real number such that $1/2\leq\beta<1$ and $L(\sigma,\chi)\neq 0$ for all $\sigma>\beta$ and all Dirichlet character $\chi$ modulo $q$.
  1. (1)
    Suppose that there exists $x_{0}\geq 2$ such that either
    $\sum_{{n\leq x}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}-\frac{4\sqrt{x}}{\varphi(q)}$ (1.21)
    has a constant sign for all $x\geq x_{0}$, or
    $\sum_{{n\leq xe^{2}}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}$ (1.22)
    has a constant sign for all $x\geq x_{0}$. Then $L(s,\chi)\not=0$ in the right half-plane $\Re(s)>\beta$ for all $\chi$ modulo $q$. In particular, if $\beta=1/2$, the GRH for $L(s,\chi)$ holds for all Dirichlet characters $\chi$ modulo $q$.
  2. (2)
    Suppose that $L(1/2,\chi)\not=0$ for all Dirichlet character $\chi$ modulo $q$. Then, we have
    $\lim_{x\to\infty}$ $\left(\sum_{{n\leq x}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)-\frac{4\sqrt{x}}{\varphi(q)\log x}\right)$ (1.23)
    $\qquad\qquad\qquad\qquad\qquad\quad=-\frac{1}{\varphi(q)}\sum_{\chi\,{\rm mod}\,q}\frac{L^{\prime}}{L}\left(\frac{1}{2},\chi\right)$
    if and only if the GRH for $L(s,\chi)$ holds for all Dirichlet characters $\chi$ modulo $q$.
  3. (3)
    Suppose that $L(1/2,\chi)\not=0$ for all Dirichlet character $\chi$ modulo $q$. Then, we have
    $\lim_{x\to\infty}\sum_{{n\leq xe^{2}}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)=-\frac{1}{\varphi(q)}\sum_{\chi\,{\rm mod}\,q}\frac{L^{\prime}}{L}\left(\frac{1}{2},\chi\right)$ (1.24)
    if and only if the GRH for $L(s,\chi)$ holds for all Dirichlet characters $\chi$ modulo $q$ and
    $\sum_{\chi\,{\rm mod}\,q}\sum_{\rho_{\chi^{\ast}}}\frac{x^{\rho_{\chi^{\ast}}}}{\rho_{\chi^{\ast}}}=o(\sqrt{x}\log x)$ (1.25)
    holds as $x\to\infty$ for nontrivial zeros of $L(s,\chi^{\ast})$ for all $\chi$ modulo $q$, where $\chi^{\ast}$ is a primitive Dirichlet character that induces $\chi$. As well as the sum $\sum_{\rho}$ above, the sum $\sum_{\rho_{\chi^{\ast}}}$ runs over all nontrivial zeros of $L(s,\chi^{\ast})$ counted with multiplicity, and is defined as $\sum_{\rho_{\chi^{\ast}}}=\lim_{T\to\infty}\sum_{|\Im(\rho_{\chi^{\ast}})|\leq T}$.
  4. (4)
    Let $m_{\chi}$ denote the order of the zero of $L(s,\chi)$ at $s=1/2$. Then, we have
    $\lim_{x\to\infty}$ $\left(\frac{1}{\log x}\sum_{{n\leq x}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)-\frac{4\sqrt{x}}{\varphi(q)(\log x)^{2}}\right)$ (1.26)
    $\qquad\qquad\qquad\qquad\qquad\quad=-\frac{1}{2}\sum_{\chi\,{\rm mod}\,q}m_{\chi}$
    if and only if the GRH for $L(s,\chi)$ holds for all Dirichlet characters $\chi$ modulo $q$.
  5. (5)
    Under the notation of the third and fourth statements, we have
    $\lim_{x\to\infty}\,\frac{1}{\log x}\sum_{{n\leq xe^{2}}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)=-\frac{1}{2}\sum_{\chi\,{\rm mod}\,q}m_{\chi}$ (1.27)
    if and only if the GRH for $L(s,\chi)$ holds for all Dirichlet characters $\chi$ modulo $q$ and
    $\sum_{\chi\,{\rm mod}\,q}\sum_{\rho_{\chi^{\ast}}}\frac{x^{\rho_{\chi^{\ast}}}}{\rho_{\chi^{\ast}}}=o(\sqrt{x}(\log x)^{2})$ (1.28)
    holds as $x\to\infty$ for nontrivial zeros of $L(s,\chi^{\ast})$ for all $\chi$ modulo $q$.
If $L(1/2,\chi)\not=0$ for all Dirichlet character $\chi$ modulo $q$, then the right-hand sides of (1.26) and (1.27) vanish. Hence, the second and fourth claims of Theorem 6, as well as the third and fifth claims, are consistent.
As for (1.10), nonpositivity of (1.22) can be expressed as an inequality for nonnegative sums
$\sum_{{n\leq x}\atop{n\equiv 1\,{\rm mod}\,q}}\Lambda(n)\sqrt{\frac{x}{n}}\log\frac{x}{n}\quad\leq~\sum_{{x<n\leq xe^{2}}\atop{n\equiv 1\,{\rm mod}\,q}}\Lambda(n)\sqrt{\frac{x}{n}}\log\frac{n}{x}.$
The previous (1.13) corresponds to the case $q=1$ of this inequality.
According to the second and third statements of Theorem 6, the sign of the sums (1.21) and (1.22) for sufficiently large $x$, which are assumed in the first statement, should be determined by the constant
$C(q):=\sum_{\chi\,{\rm mod}\,q}\frac{L^{\prime}}{L}\left(\frac{1}{2},\chi\right)\in\mathbb{R}.$
For simplicity, let $q$ be an odd prime. Then, taking into account that there are (q1)/21(q-1)/2-1 even primitive Dirichlet characters and (q1)/2(q-1)/2 odd primitive Dirichlet characters, we obtain the formula
$C(q)=\frac{q-1}{2}(\log 8\pi+C_{0})-\frac{q-2}{2}\log q+\frac{\log q}{\sqrt{q}-1}$
from (3.2), (4.11), and (4.12) below, where $C_{0}$ is the Euler–Mascheroni constant. It follows that $C(q)<0$ for all odd primes $\leq 47$, while $C(q)>0$ for all primes $\geq 53$.
The sign constancy of (1.22) is conditional, however, we find that (1.22) is unconditionally negative on average for all sufficiently large $x>1$.
Theorem 7.
We have
$\sum_{3\leq q\leq Q}$ $\left(\sum_{\chi\,{\rm mod}\,q}\sum_{n\leq x}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}\log\frac{x}{n}-4\sqrt{x}\right)$ (1.29)
$\qquad\qquad\qquad\qquad=4\sqrt{x}\,\left[\frac{1}{9}x\,(1+o(1))-Q\right]\quad\text{as $x\to\infty$}$
and
$\sum_{3\leq q\leq Q}\sum_{\chi\,{\rm mod}\,q}\sum_{n\leq xe^{2}}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}\log\frac{x}{n}=-\frac{8}{9}\,e^{3}\,x\sqrt{x}\,(1+o(1))\quad\text{as $x\to\infty$}$ (1.30)
if $Q\geq x$. In particular, the left-hand sides are negative for all sufficiently large $x>1$.
The reason $q=1,2$ are excluded in both (1.29) and (1.30) is that, in those cases, the Dirichlet characters modulo $q$ are only the principal character. While it is certainly desirable to obtain the negativity of the left-hand side of (1.29) or (1.30) for a large constant $Q\geq 3$, we are also interested in what nontrivial conclusions can be derived when the negativity of the left-hand side of (1.29) or (1.30) is obtained under a condition weaker than $Q\geq x$ with respect to $Q$. This remains a subject for future study.
This paper is organized as follows. In Section 2, we review the basic properties of the Riemann zeta function and Dirichlet $L$-functions. In Section 3, we prove the results for the Riemann zeta function: Theorems 1 and 2 and Corollary 1. The key ingredients for the proof of Theorem 1 are an integral representation for a function involving the logarithmic derivative of the Riemann zeta function and a well-known result concerning the Laplace transform of a nonnegative valued function. Except for technical details, all results in this paper are shown using roughly the same approach as Theorem 1. In Section 4, we prove the results for Dirichlet $L$-functions: Theorems 3 and 4. Although Theorem 5 is a result for the Riemann zeta function, it follows a similar way to that of Theorem 4, so we include its proof in this section. Finally, in Section 5, we prove Theorems 6 and 7 by combining the discussions in Sections 3 and 4.
Acknowledgments  The author appreciates the referee’s valuable comments, which led to the corrections of Theorems 6, 8, and 9, and contributed to improving the presentation of the paper. This work was supported by JSPS KAKENHI Grant Number JP23K03050.
2. Preliminaries
In this section, we review a few elementary facts in number theory and the basic properties of the Riemann zeta function and Dirichlet $L$-functions referencing Titchmarsh [23] and Montgomery and Vaughan [15]. In addition, we review a property of the Mellin transforms of nonnegative functions.
2.1. Facts on prime numbers
We use Mertens’ theorem [15, Theorem 2.7] in the form
$\sum_{p\leq x}\frac{\log p}{p}=\log x+O(1)\quad\text{as $x\to\infty$},$ (2.1)
and the classical prime number theorem [15, Theorem 6.9]
$\psi(x)-x=\sum_{\rho}\frac{x^{\rho}}{\rho}=O(x\exp(-c\sqrt{\log x}))\quad\text{as $x\to\infty$},$ (2.2)
where $\sum_{\rho}=\lim_{T\to\infty}\sum_{|\Im(\rho)|\leq T}$ as before.
2.2. Riemann zeta function
The Riemann zeta function $\zeta(s)$ is defined by the absolutely convergent series $\sum_{n=1}^{\infty}n^{-s}$ if $\Re(s)>1$, and extends to a meromorphic function on $\mathbb{C}$, which is analytic except for a simple pole at $s=1$ with residue one. Due to the Euler product formula $\zeta(s)=\prod_{p}(1-p^{-s})^{-1}$ [23, (1.1.2)], the logarithmic derivative of $\zeta(s)$ has the Dirichlet series
$\frac{\zeta^{\prime}}{\zeta}(s)=\sum_{n=1}^{\infty}\frac{\Lambda(n)}{n^{s}}$ (2.3)
for $\Re(s)>1$ [23, (1.1.8)]. The Riemann xi-function (1.8) is an entire function satisfying the functional equations
$\xi(s)=\xi(1-s)\quad\text{and}\quad\xi(s)=\overline{\xi(\bar{s})}$ (2.4)
by [23, (2.6.4)] and $\zeta(\sigma)\in\mathbb{R}$ for $\sigma>1$. Zeros of $\zeta(s)$ in the critical strip $0\leq\Re(s)\leq 1$ are called nontrivial zeros and coincide with zeros of $\xi(s)$. Because $\xi(s)$ is an entire function of order one [23, Theorem 2.12],
$\sum_{\rho}|\rho|^{-1-\delta}<\infty\quad\text{if $\delta>0$},$ (2.5)
where $\rho$ runs over all nontrivial zeros of $\zeta(s)$ counted with multiplicity. We often use the explicit formula
$\sideset{}{{}^{\prime}}{\sum}_{n\leq x}\frac{\Lambda(n)}{n^{s}}$ $=\frac{x^{1-s}}{1-s}-\frac{\zeta^{\prime}}{\zeta}(s)-\sum_{\rho}\frac{x^{\rho-s}}{\rho-s}+\sum_{k=1}^{\infty}\frac{x^{-2k-s}}{2k+s}$ (2.6)
for $x>1$, $s\not=1,\rho,-2k$ ($k\in\mathbb{Z}_{>0}$) [15, Exercises 12.1.1, 4].
2.3. Dirichlet $L$-function
For a Dirichlet character $\chi$ modulo $q>1$, the Dirichlet $L$-function $L(s,\chi)$ is defined by the absolutely convergent series $\sum_{n=1}^{\infty}\chi(n)n^{-s}$ if $\Re(s)>1$. If $\chi$ is a principal character $\chi_{0}$, then $L(s,\chi_{0})$ extends to a meromorphic function on $\mathbb{C}$, which is analytic except for a simple pole at $s=1$. If $\chi$ is not a principal character, then $L(s,\chi)$ extends to an entire function and
$L(1,\chi)\not=0$ (2.7)
[15, Theorem 4.9]. Due to the Euler product formula
$L(s,\chi)=\prod_{p}(1-\chi(p)p^{-s})^{-1}\quad$ (2.8)
for $\Re(s)>1$ [15, (4.21)], the logarithmic derivative of $L(s,\chi)$ has the absolutely convergent Dirichlet series
$\frac{L^{\prime}}{L}(s,\chi)=\sum_{n=1}^{\infty}\frac{\Lambda(n)\chi(n)}{n^{s}}$ (2.9)
for $\Re(s)>1$ [15, (4.25)].
For a primitive Dirichlet character $\chi$ modulo $q>1$, we set
$\kappa:=\kappa(\chi)=\begin{cases}~0&\text{if $\chi(-1)=1$},\\ ~1&\text{if $\chi(-1)=-1$},\end{cases}$
and
$\xi(s,\chi):=\left(\frac{q}{\pi}\right)^{(s+\kappa)/2}\Gamma\left(\frac{s+\kappa}{2}\right)L(s,\chi).$ (2.10)
Then, it is an entire function satisfying the functional equation
$\xi(s,\chi)=\varepsilon(\chi)\xi(1-s,\overline{\chi}),\quad\varepsilon(\chi):=\frac{\tau(\chi)}{i^{\kappa}\sqrt{q}}$ (2.11)
[15, (10.17) and Corollary 10.8], where $\tau(\chi)$ is the Gauss sum of $\chi$. Zeros of $L(s,\chi)$ in the critical strip $0\leq\Re(s)\leq 1$ and $s\not=0$ are called nontrivial zeros and coincide with the zeros of $\xi(s,\chi)$. The Generalized Riemann Hypothesis (GRH) claims that all nontrivial zeros lie on the critical line $\Re(s)=1/2$. Because $\xi(s,\chi)$ is an entire function of order one by [15, (10.31)],
$\sum_{\rho_{\chi}}|\rho_{\chi}|^{-1-\delta}<\infty\quad\text{if $\delta>0$},$ (2.12)
where $\rho_{\chi}$ runs over all nontrivial zeros of $L(s,\chi)$ counted with multiplicity. By an argument similar to the proof of (2.6), we obtain
$\sideset{}{{}^{\prime}}{\sum}_{n\leq x}\frac{\Lambda(n)\chi(n)}{n^{s}}=-\frac{L^{\prime}}{L}(s,\chi)-\sum_{\rho_{\chi}}\frac{x^{\rho_{\chi}-s}}{\rho_{\chi}-s}+\sum_{k=0}^{\infty}\frac{x^{-2k-\kappa-s}}{2k+\kappa+s}$ (2.13)
for $x>1$, $s\not=\rho_{\chi},-2k-\kappa$ ($k\in\mathbb{Z}_{\geq 0}$) [26, (6)].
Suppose that $L(s,\chi)$ has a zero of order $m\geq 1$ at $s=1/2$. By isolating the term of $\rho_{\chi}=1/2$ in the second sum on the right-hand side of (2.13), then taking the limit $s\to 1/2$, we obtain
$\sideset{}{{}^{\prime}}{\sum}_{n\leq x}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}$ $=-m\log x-\lim_{s\to 1/2}\left[\frac{L^{\prime}}{L}(s,\chi)-\frac{m}{s-1/2}\right]$ (2.14)
$\quad-\sum_{\rho_{\chi}\not=1/2}\frac{x^{\rho_{\chi}-1/2}}{\rho_{\chi}-1/2}+\sum_{k=0}^{\infty}\frac{x^{-2k-\kappa-1/2}}{2k+\kappa+1/2}.$
If $\chi$ is imprimitive, it is induced from a primitive Dirichlet character $\chi^{\ast}$ modulo $q^{\ast}$, for some $q^{\ast}\mid q$, that is,
$\chi(n)=\begin{cases}~\chi^{\ast}(n)&(n,q)=1,\\ ~0&\text{otherwise}.\end{cases}$
Therefore,
$L(s,\chi)=L(s,\chi^{\ast})\prod_{p\mid q}(1-\chi^{\ast}(p)p^{-s})$ (2.15)
[15, (9.1), (9.2)]. For a principal character $\chi_{0}$ modulo $q$, $L(s,\chi_{0})=\zeta(s)\prod_{p\mid q}(1-p^{-s})$. We define the GRH to hold for $L(s,\chi)$ as the GRH to hold for $L(s,\chi^{\ast})$.
2.4. Mellin transforms of nonnegative functions
We state a well-known property of the Laplace transform of nonnegative functions in terms of the Mellin transform [25, Theorem 5b in Chap. II] (or [15, Lemma 15.1]). This is an analog of Landau’s Theorem concerning Dirichlet series with nonnegative coefficients.
Proposition 1.
If $f(x)$ is nonnegative, then the real point of the axis of convergence of
$F(s)=\int_{1}^{\infty}f(x)\,x^{-s}\,\frac{dx}{x}$
is a singularity of $F(s)$.
3. Proof of Theorems 1 and 2
3.1. Proof of Theorem 1
The equality in (1.5) is obtained by integrating (1.3). First, we prove (1.5) and (1.6) assuming the RH. We set
$f(x):=\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}.$
By applying [20, Lemma 1] to $s=1/2$, we obtain
$f(x)-4\sqrt{x}$ $=-\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)\log x-\left(\frac{\zeta^{\prime}}{\zeta}\right)^{\prime}\left(\frac{1}{2}\right)$ (3.1)
$\quad-\sum_{\rho}\frac{x^{\rho-1/2}}{(\rho-1/2)^{2}}-\sum_{k=1}^{\infty}\frac{x^{-2k-1/2}}{(2k+1/2)^{2}}$
for $x\geq 2$. The second sum on the second line of the right-hand side is clearly bounded for $x\geq 2$. The first sum on the second line of the right-hand side is also bounded for $x\geq 2$, because $\rho-1/2$ is purely imaginary by the RH and (2.5). On the other hand,
$\frac{\xi^{\prime}}{\xi}\left(\frac{1}{2}\right)=\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)+\frac{1}{2}\left[\frac{\Gamma^{\prime}}{\Gamma}\left(\frac{1}{4}\right)-\log\pi\right]=0$
by the taking logarithmic derivative of the first equation in (2.4) and using $(\Gamma^{\prime}/\Gamma)(1/4)=-\pi/2-3\log 2-C_{0}=-4.22745\dots<0$ ([15, p. 411]). Therefore,
$\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)=\frac{1}{2}\left(\frac{\pi}{2}+\log 8\pi+C_{0}\right)=2.68609\dots>0.$ (3.2)
Hence, (1.5) holds for all sufficiently large $x\geq 2$. Furthermore, dividing both sides of (3.1) by $\log x$ and taking the limit as $x\to\infty$, we obtain (1.6).
By (3.2), equality (1.6) implies that (1.5) holds for all sufficiently large $x$. Therefore, it remains to show that the RH holds if there exist $x_{0}\geq 2$ such that (1.5) is valid for all $x\geq x_{0}$. The key point is the integral formula
$\int_{1}^{\infty}(-f(x))\,x^{-s+1/2}\,\frac{dx}{x}=\frac{1}{(s-1/2)^{2}}\frac{\zeta^{\prime}}{\zeta}(s)\quad\text{for}~\Re(s)>1,$ (3.3)
which is derived as follows. For an interval $I\subset\mathbb{R}$, we denote by $\mathbf{1}_{I}$ the indicator function of $I$. Using the integral formula
$\int_{1}^{\infty}\frac{1}{\sqrt{n}}\log\frac{x}{n}\cdot\mathbf{1}_{[n,\infty)}(x)\cdot x^{-s+1/2}\,\frac{dx}{x}=\frac{1}{(s-1/2)^{2}}\,n^{-s}\quad\text{for}~\Re(s)>1/2$
obtained by a straightforward calculation for any $n\in\mathbb{Z}_{>0}$, we have
$\int_{1}^{\infty}(-f(x))\,x^{-s+1/2}\,\frac{dx}{x}$ $=-\sum_{n=2}^{\infty}\Lambda(n)\int_{1}^{\infty}\frac{1}{\sqrt{n}}\log\frac{x}{n}\cdot\mathbf{1}_{[n,\infty)}(x)\cdot x^{-s+1/2}\,\frac{dx}{x}$
$=\frac{1}{(s-1/2)^{2}}\left(-\sum_{n=2}^{\infty}\Lambda(n)\,n^{-s}\right)$
if $\Re(s)>1$, where the interchange of the order of integration and summation is justified by the absolute convergence of the series for $\Re(s)>1$ and Fubini’s theorem. Therefore, we obtain (3.3) by (2.3).
On the other hand, $(\zeta^{\prime}/\zeta)(s)$ has no poles in the half-line $(1/2,\infty)$ on the real axis except for the simple pole at $s=1$ with residue 1-1, because $\zeta(s)$ has the simple pole at $s=1$ with residue one and has no zeros in the half-line $(1/2,\infty)$ by the series expansion $(1-2^{1-s})\zeta(s)=\sum_{n=1}^{\infty}(-1)^{n-1}n^{-s}>0$ for $s>0$.
By the integral formula (3.3) and $\int_{1}^{\infty}\sqrt{x}\cdot x^{-s+1/2}dx/x=1/(s-1)$ for $\Re(s)>1$,
$\int_{1}^{\infty}$ $(-f(x)+4\sqrt{x}\,)x^{-s+1/2}\,\frac{dx}{x}$ (3.4)
$=\frac{1}{(s-1/2)^{2}}\frac{\zeta^{\prime}}{\zeta}(s)+\frac{4}{s-1}\quad\text{for}~\Re(s)>1.$
The simple pole of $(s-1/2)^{-2}(\zeta^{\prime}/\zeta)(s)$ at $s=1$ is canceled out by the second term of the right-hand side. If we divide the left-hand side of (3.4) into
$\left(\int_{x_{0}}^{\infty}+\int_{1}^{x_{0}}\right)(-f(x)+4\sqrt{x}\,)x^{-s+1/2}\,\frac{dx}{x},$
the second integral defines an entire function. Applying Proposition 1 to the first integral, we find that the abscissa of convergence of this integral is at most $1/2$. Therefore, the integral converges uniformly on any compact subset of the right half-plane $\Re(s)>1/2$, and hence defines an analytic function there. This implies, by (3.4), that $(\zeta^{\prime}/\zeta)(s)$ has no poles in the right half-plane $\Re(s)>1/2$, which in turn implies that the RH holds. $\Box$
3.2. Proof of Corollary 1
Assuming the RH, it is shown that $g_{0}(t)$ is nonnegative for all sufficiently large $t>0$ by (1.9) and (3.1) as in the proof of Theorem 1. Therefore, we prove the converse. Using (1.9), (3.3), and the integral formula
$\int_{0}^{\infty}4(e^{t/2}+e^{-t/2}-2)\,e^{-t(s-1/2)}\,dt=\frac{1}{(s-1/2)^{2}}\,\left(\frac{1}{s-1}+\frac{1}{s}\right)\quad\text{for}~\Re(s)>1,$
we have
$\int_{0}^{\infty}(-g_{0}(t))\,e^{-t(s-1/2)}\,dt=\frac{1}{(s-1/2)^{2}}\left(\frac{\zeta^{\prime}}{\zeta}(s)+\frac{1}{s-1}+\frac{1}{s}\right)$ (3.5)
for $\Re(s)>1$. The simple pole of $(\zeta^{\prime}/\zeta)(s)$ at $s=1$ is canceled out by the second term of the right-hand side. Hence, if we use (3.5) instead of (3.4), the RH follows from the nonnegativity of $g_{0}(t)$ for all sufficiently large $t>0$ as in the proof of Theorem 1. $\Box$
3.3. Proof of Theorem 2
First, we prove that the nonpositivity (1.10) for all sufficiently large $x>0$ implies the RH. Using the integral formula
$\int_{1}^{\infty}\frac{1}{\sqrt{n}}\cdot\mathbf{1}_{[n,\infty)}(x)\cdot x^{-s+1/2}\,\frac{dx}{x}=\frac{1}{s-1/2}\,n^{-s}\quad\text{for}~\Re(s)>1/2$
obtained by a straightforward calculation for any $n\in\mathbb{Z}_{>0}$, we have
$\int_{1}^{\infty}\left(\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\right)x^{-s+1/2}\,\frac{dx}{x}$ $=-\frac{1}{s-1/2}\frac{\zeta^{\prime}}{\zeta}(s)$
for $\Re(s)>1$. Combining this with (3.3) and denoting $2$ as $\log\exp(2)$, we obtain
$\int_{1}^{\infty}\left(-\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{ne^{2}}\right)x^{-s+1/2}\,\frac{dx}{x}=-\frac{2(s-1)}{(s-1/2)^{2}}\frac{\zeta^{\prime}}{\zeta}(s)\quad\text{for}~\Re(s)>1.$ (3.6)
On the right-hand side, as stated in the proof of Theorem 1, $(\zeta^{\prime}/\zeta)(s)$ has no poles in the half-line $(1/2,\infty)$ except for the simple pole at $s=1$, but it is canceled out by s1s-1. Hence, it is shown that the nonpositivity (1.10) for sufficiently large $x>0$ implies the RH as in the proof of Theorem 1.
We prove the second half of the theorem. Applying (2.6) to $s=1/2$,
$2\sqrt{x}=\psi_{1/2}(x)+\sum_{\rho}\frac{x^{\rho-1/2}}{\rho-1/2}+\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)-\sum_{k=1}^{\infty}\frac{x^{-2k-1/2}}{2k+1/2}$ (3.7)
for $x>1$. Substituting (3.7) into (3.1),
$\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}-2\sideset{}{{}^{\prime}}{\sum}_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}$ $=-\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)\log x+2\sum_{\rho}\frac{x^{\rho-1/2}}{\rho-1/2}$ (3.8)
$\quad-\sum_{\rho}\frac{x^{\rho-1/2}}{(\rho-1/2)^{2}}-\left(\frac{\zeta^{\prime}}{\zeta}\right)^{\prime}\left(\frac{1}{2}\right)+2\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)$
$\quad-\sum_{k=1}^{\infty}\frac{x^{-2k-1/2}}{(2k+1/2)^{2}}-2\sum_{k=1}^{\infty}\frac{x^{-2k-1/2}}{2k+1/2}.$
By denoting $2$ as $\log\exp(2)$, the left-hand side of (3.8) is written as
$\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{ne^{2}}+O\left(\frac{1}{\sqrt{x}}\log x\right),$
where the error term arises by replacing $\sideset{}{{}^{\prime}}{\sum}_{n\leq x}$ with $\sum_{n\leq x}$. Substituting this into (3.8), and then replacing $x$ with $xe^{2}$ and dividing by $\log x$, we have
$\sum_{n\leq xe^{2}}\frac{\Lambda(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)$ $=-\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)+\frac{2}{\log x}\sum_{\rho}\frac{(xe^{2})^{\rho-1/2}}{\rho-1/2}$ (3.9)
$\quad-\frac{1}{\log x}\sum_{\rho}\frac{(xe^{2})^{\rho-1/2}}{(\rho-1/2)^{2}}+O\left(\frac{1}{\log x}\right)$
as $x\to\infty$.
Suppose that equation (1.11) holds. Then, inequality (1.10) is valid for all sufficiently large $x>0$ by (3.2), and therefore, the RH holds from the first half of the theorem. From the RH, the third term on the right hand side of (3.9) is $O(1/\log x)$, and
$\sum_{\rho}\frac{x^{\rho-1/2}}{\rho-1/2}-\sum_{\rho}\frac{x^{\rho-1/2}}{\rho}=\frac{1}{2}\sum_{\rho}\frac{x^{\rho-1/2}}{\rho(\rho-1/2)}=O(1).$
Hence, (1.11) implies that (1.12) holds. Conversely, assuming that the RH and (1.12) hold, (1.11) follows from (3.9). $\Box$
4. Proof of Theorems 3, 4, and 5
4.1. Proof of Theorem 3
We prove the following theorem as Theorem 3 is obtained by applying it to $\chi=\chi_{4}$.
Theorem 8.
Let $\chi$ be a non-principal Dirichlet character of modulo $q$, and denote $L(s,\chi)$ by $L(s)$. Suppose that there exists $x_{0}\geq 2$ such that
$\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\Re(\chi(n))\log\frac{x}{n}$ (4.1)
has a constant sign for all $x\geq x_{0}$, and $L(\sigma)\not=0$ for $\sigma>\beta$ for some $1/2\leq\beta<1$. Then, $L(s)\not=0$ in the right-half plane $\Re(s)>\beta$. In particular, if (4.1) has a constant sign for all sufficiently large $x>0$, and $L(\sigma)\not=0$ for $\sigma>1/2$, then the GRH for $L(s)$ holds.
Suppose that $L(1/2)\not=0$. Then, we have
$\lim_{x\to\infty}\sum_{n\leq x}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)=-\frac{L^{\prime}}{L}\left(\frac{1}{2}\right)$ (4.2)
if and only if the GRH for $L(s,\chi)$ holds. If $\chi$ is primitive, then equation (4.2) implies that the sum (4.1) has a constant sign for all sufficiently large $x>0$.
Furthermore, suppose that $s=1/2$ is a zero of $L(s)$ of order $m\geq 0$. Then, we have
$\lim_{x\to\infty}\frac{1}{\log x}\sum_{n\leq x}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)=-\frac{m}{2}$ (4.3)
if and only if the GRH for $L(s,\chi)$ holds. Equation (4.3) implies that the value of the sum (4.1) is negative for all sufficiently large $x>0$ if $m\geq 1$.
Remark. It is conjectured that $L(\sigma,\chi)\not=0$ for all $1/2\leq\sigma\leq 1$ and for all $\chi$. This conjecture is confirmed for primitive Dirichlet characters of modulus $q\leq 4\cdot 10^{5}$ by Platt [17], and for real odd Dirichlet characters of modulus $q\leq 3\cdot 10^{8}$ by Watkins [24].
Proof.
We provide only a sketch of the proof, as the argument is similar to that of Theorems 1 and 2. However, we carefully address the points that require particular consideration in the case of $L(s)$. We set
$f_{\chi}(x):=\sum_{n\leq x}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}\log\frac{x}{n}.$ (4.4)
Then, the integral formula
$\int_{1}^{\infty}(-f_{\chi}(x))\,x^{-s+1/2}\,\frac{dx}{x}=\frac{1}{(s-1/2)^{2}}\frac{L^{\prime}}{L}(s)\quad\text{for}~\Re(s)>1$ (4.5)
is obtained in the same way as (3.3) using the series expansion (2.9). On the right-hand side, there exists $1/2\leq\beta<1$ such that $L(\sigma)\not=0$ for any real number $\sigma>\beta$, because $L(s)$ is analytic in $\mathbb{C}$, and has no poles and zeros in the half-line $1,\infty)\subset\mathbb{R}$ by the Euler product (2.8) and Dirichlet’s theorem (2.7). Therefore, if $\chi$ is real, the first half of the theorem follows similarly to the proof of Theorem 1.
We consider the case where $\chi$ is not real. We replace $\chi$ by $\bar{\chi}$ in (4.5). Combining this with (4.5), we have
$\int_{1}^{\infty}\left(-\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\Re(\chi(n))\log\frac{x}{n}\right)$ $x^{-s+1/2}\,\frac{dx}{x}$ (4.6)
$=\frac{1}{(s-1/2)^{2}}\left(\frac{L^{\prime}}{L}(s,\chi)+\frac{L^{\prime}}{L}(s,\bar{\chi})\right).$
By $L(s,\bar{\chi})=\overline{L(\bar{s},\chi)}$, the assumption that $L(\sigma,\chi)\neq 0$ for $\sigma>\beta$ implies $L(\sigma,\bar{\chi})\neq 0$ for $\sigma>\beta$. This yields that the right-hand side of (4.6) has no poles on the interval $(\beta,\infty)\subset\mathbb{R}$. Thus, the assumption on (4.1) implies that the abscissa of convergence for the integral on the left-hand side of (4.6) is at most $\beta$ thanks to Proposition 1. This means that the right-hand side of (4.6) is analytic in $\Re(s)>\beta$. This forces $L(s,\chi)\neq 0$ in $\Re(s)>\beta$, since the residues of the logarithmic derivative at the zeros are positive integers and hence cannot cancel between $(L^{\prime}/L)(s,\chi)$ and $(L^{\prime}/L)(s,\bar{\chi})$.
We prove the second half of the theorem. Let $\chi^{\ast}$ be a primitive character that induces $\chi$. Then, by applying the argument in the proof of [20, Lemma 1] to $L(s)$, we obtain
$\sum_{n\leq x}\frac{\Lambda(n)\chi(n)}{n^{s}}\log\frac{x}{n}$ $=-\frac{L^{\prime}}{L}(s)\log x-\sum_{\rho}\frac{x^{\rho-s}}{(\rho-s)^{2}}$ (4.7)
$\quad-\left(\frac{L^{\prime}}{L}\right)^{\prime}(s)-\sum_{k=0}^{\infty}\frac{x^{-2k-\kappa^{\ast}-s}}{(2k+\kappa^{\ast}+s)^{2}}$
for $x>1$, $s\not=\rho_{\chi},-2k-\kappa$ ($k\in\mathbb{Z}_{\geq 0}$). If $L(1/2)\not=0$, that is, if $m=0$, then substituting $s=1/2$ into (4.7) yields
$f_{\chi}(x)$ $=-\frac{L^{\prime}}{L}\left(\frac{1}{2}\right)\log x-\sum_{\rho}\frac{x^{\rho-1/2}}{(\rho-1/2)^{2}}$ (4.8)
$\quad-\left(\frac{L^{\prime}}{L}\right)^{\prime}\left(\frac{1}{2}\right)-\sum_{k=0}^{\infty}\frac{x^{-2k-\kappa^{\ast}-1/2}}{(2k+\kappa^{\ast}+1/2)^{2}},$
where the sum $\sum_{\rho}$ is taken over all nontrivial zeros of $L(s,\chi^{\ast})$, which coincide with all zeros of $\xi(s,\chi^{\ast})$, and the zeros of the finite product $L(s)/L(s,\chi^{\ast})=\prod_{p\mid q}(1-\chi^{\ast}(p)p^{-s})$, and $\kappa^{\ast}=\kappa(\chi^{\ast})$.
Suppose that $L(s)$ has a zero of order $m\geq 1$ at $s=1/2$. Then, we have
$\lim_{s\to 1/2}$ $\left(-\frac{L^{\prime}}{L}(s)\log x-\frac{x^{-s+1/2}}{(-s+1/2)^{2}}-\left(\frac{L^{\prime}}{L}\right)^{\prime}(s)\right)$ (4.9)
$=-\frac{m}{2}(\log x)^{2}-\lim_{s\to 1/2}\left[\frac{L^{\prime}}{L}(s,\chi)-\frac{m}{s-1/2}\right]\,\log x$
$\quad-\lim_{s\to 1/2}\left[\left(\frac{L^{\prime}}{L}\right)^{\prime}(s)+\frac{m}{(s-1/2)^{2}}\right].$
Therefore, by isolating the term of $\rho=1/2$ in the second sum on the right-hand side of (4.7), then taking the limit $s\to 1/2$, we obtain
$f_{\chi}(x)$ $=-\frac{m}{2}(\log x)^{2}-\lim_{s\to 1/2}\left[\frac{L^{\prime}}{L}(s,\chi)-\frac{m}{s-1/2}\right]\,\log x$ (4.10)
$\quad-\lim_{s\to 1/2}\left[\left(\frac{L^{\prime}}{L}\right)^{\prime}(s)+\frac{m}{(s-1/2)^{2}}\right]$
$\quad-\sum_{\rho\not=1/2}\frac{x^{\rho-1/2}}{(\rho-1/2)^{2}}-\sum_{k=0}^{\infty}\frac{x^{-2k-\kappa^{\ast}-1/2}}{(2k+\kappa^{\ast}+1/2)^{2}},$
Assuming the GRH for $L(s)$, each term of the right-hand side (4.8) is bounded except for the first term, and each term of the right-hand side (4.10) is bounded except for the first and second terms. Hence, (4.2) and (4.3) holds.
Conversely, assuming (4.2) (resp. (4.3)), we have $f_{\chi}(x)\ll\log x$ (resp. $\ll(\log x)^{2}$) as $x\to\infty$. Therefore, the integral on the left-hand side of (4.5) converges uniformly on any compact subset in $\Re(s)>1/2$. Thus, it defines an analytic function in $\Re(s)>1/2$, and hence, $L(s)$ has no zeros (or poles) in $\Re(s)>1/2$, that is the GRH for $L(s)$.
Finally, we prove that the sum (4.1) has a constant sign for all sufficiently large $x>0$ assuming $L(1/2)\neq 0$, (4.2) and that $\chi$ is primitive. Taking the logarithmic derivative of (2.11) and substituting $s=1/2$, we obtain $\Re(\xi^{\prime}/\xi)(1/2,\chi)=0$. This gives
$\Re\frac{L^{\prime}}{L}\left(\frac{1}{2}\right)=\frac{1}{2}\left[\log\frac{\pi}{q}-\frac{\Gamma^{\prime}}{\Gamma}\left(\frac{1}{4}+\frac{\kappa(\chi)}{2}\right)\right]$ (4.11)
by (2.10). In particular, $\Re(L^{\prime}/L)(1/2)$ depends only on $q$ and the parity of $\chi$. By (4.11), the value $\Re(L^{\prime}/L)(1/2)$ vanishes only when $q>1$ takes the following values
$q=\pi\cdot\exp\left[-\frac{\Gamma^{\prime}}{\Gamma}\left(\frac{1}{4}+\frac{\kappa(\chi)}{2}\right)\right]=\begin{cases}~215.332\cdots,&\kappa=0,\\ ~9.305\cdots,&\kappa=1,\end{cases}$
however it is impossible for any integer $q>1$, where
$\frac{\Gamma^{\prime}}{\Gamma}\left(\frac{1}{4}+\frac{\kappa(\chi)}{2}\right)=\left(\kappa(\chi)-\frac{1}{2}\right)\pi-3\log 2-C_{0}$ (4.12)
([15, p. 411]). Therefore, $\Re(L^{\prime}/L)(1/2)$ is a nonzero real number. Hence, by taking the real part of (4.2), it follows that (4.1) has a constant sign for all sufficiently large $x>0$. ∎
Proof of Theorem 3. From the second half of Theorem 8, we see that the first and third claims of Theorem 3 are equivalent. Using the series expansion $\lim_{N\to\infty}\sum_{n=0}^{N}(-1)^{n}(2n+1)^{-s}$ for $\Re(s)>0$, we find that $L(s,\chi_{4})$ has no zeros on the positive real line $(0,\infty)$. Therefore, by the first half of Theorem 8, the second claim of Theorem 3 implies the first claim. Assuming the first claim of Theorem 3, it follows that the third claim holds by the equivalence we have already shown. Furthermore, since $\chi_{4}$ is primitive, the third claim implies the second claim. $\Box$
4.2. Proof of Theorem 4
We prove the following theorem as Theorem 4 is obtained by applying it to $\chi=\chi_{4}$. Note that $L(1/2,\chi_{4})\neq 0$ as mentioned in the proof of Theorem 3.
Theorem 9.
Let $\chi$ be a real Dirichlet character such that $L(\sigma,\chi)\not=0$ for $\sigma>\beta~(\geq 1/2)$. Suppose that
$\sum_{p\leq x}\chi(p)\log p\cdot\sqrt{\frac{x}{p}}\,\log\frac{x}{p}\leq 0$ (4.13)
for all sufficiently large $x>1$. Then, $L(s,\chi)\not=0$ in the right-half plane $\Re(s)>\beta$. In particular, if $\beta=1/2$, the GRH for $L(s,\chi)$ holds.
Conversely, suppose that the GRH for $L(s,\chi)$ holds and $L(1/2,\chi)\neq 0$. Then, the asymptotic formula
$\sum_{p\leq x}\chi(p)\log p\cdot\sqrt{\frac{x}{p}}\,\log\frac{x}{p}=-\frac{\sqrt{x}}{4}(\log x)^{2}\left(1+O\left(\frac{1}{\log x}\right)\right)$ (4.14)
holds as $x\to\infty$. Therefore, (4.13) is valid for all sufficiently large $x>1$, and the left-hand side of (4.13) tends to -$\infty$ as $x\to\infty$.
Proof.
Let $f_{\chi}(x)$ be the function in (4.4). If we define
$f_{1}(x):=\sum_{p\leq x}\frac{\chi(p)\log p}{\sqrt{p}}\log\frac{x}{p},$
$f_{2}(x):=\sum_{p\leq\sqrt{x}}\frac{\chi(p)^{2}\log p}{p}\log\frac{x}{p^{2}},\qquad f_{3}(x):=\sum_{k\geq 3}\sum_{p\leq x^{1/k}}\frac{\chi(p)^{k}\log p}{p^{k/2}}\log\frac{x}{p^{k}},\\$
then $f_{\chi}(x)=f_{1}(x)+f_{2}(x)+f_{3}(x)$. For $f_{2}(x)$, we have
$f_{2}(x)$ $=\log x\sum_{p\leq\sqrt{x}}\frac{\log p}{p}-2\sum_{p\leq\sqrt{x}}\frac{(\log p)^{2}}{p}+O(\log x)$
$=\frac{1}{2}(\log x)^{2}+O(\log x)-2\sum_{p\leq\sqrt{x}}\frac{(\log p)^{2}}{p}$
by using (2.1). For the third term on the right-hand side,
$\sum_{p\leq x}\frac{(\log p)^{2}}{p}$ $=\log x\sum_{p\leq x}\frac{\log p}{p}-\int_{1}^{x}\left(\sum_{p\leq t}\frac{\log p}{p}\right)\frac{1}{t}\,dt$
$=(\log x)^{2}+O(\log x)-\int_{1}^{x}(\log t+O(1))\frac{1}{t}\,dt$
$=\frac{1}{2}(\log x)^{2}+O(\log x)$
by partial summation and (2.1). Therefore,
$f_{2}(x)=\frac{1}{4}(\log x)^{2}+O(\log x).$ (4.15)
On the other hand,
$f_{3}(x)=O(\log x)$ (4.16)
by the convergence of the sum. As a consequence, if we rewrite (4.5) as
$-\int_{1}^{\infty}f_{1}(x)\,x^{-s+1/2}\frac{dx}{x}$ $=\frac{1}{(s-1/2)^{2}}\frac{L^{\prime}}{L}(s)$ (4.17)
$\quad+\int_{1}^{\infty}f_{2}(x)\,x^{-s+1/2}\frac{dx}{x}+\int_{1}^{\infty}f_{3}(x)\,x^{-s+1/2}\frac{dx}{x},$
the second and third terms on the right-hand side converge absolutely and uniformly on any compact subset in $\Re(s)>1/2$, and hence define analytic functions there.
Suppose that $f_{1}(x)$ has a constant sign for all sufficiently large $x>1$. Then, the assumption $L(\sigma,\chi)\not=0$ for $\sigma>\beta~(\geq 1/2)$ implies that the abscissa of convergence $\sigma_{c}$ for the integral on the left-hand side of (4.17) satisfies $\sigma_{c}\leq\beta$ by Proposition 1. Therefore, the integral on the left-hand side of (4.17) converges uniformly for any compact subset of $\Re(s)>\sigma_{c}$. This means that the right-hand side of (4.17) is analytic in $\Re(s)>\beta~(\geq 1/2)$. This forces $L(s,\chi)\neq 0$ in $\Re(s)>\beta~(\geq 1/2)$.
Conversely, suppose that the GRH for $L(s)$ holds. Then, $f_{\chi}(x)=O(\log x)$ by (4.8). On the other hand, $f_{\chi}(x)=f_{1}(z)+(1/4)(\log x)^{2}+O(\log x)$ by (4.15) and (4.16). Combining these two,
$f_{1}(x)=-\frac{1}{4}(\log x)^{2}\left(1+O\left(\frac{1}{\log x}\right)\right).$
This immediately implies (4.14). ∎
4.3. Proof of Theorem 5
It suffices to show (1.20) with $x$ replaced by $x/e^{2}$. Using the decomposition
$\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{ne^{2}}$ $=\sum_{p\leq x}\frac{\log p}{\sqrt{p}}\left(\log\frac{x}{p}-2\right)$
$\quad+\sum_{p\leq\sqrt{x}}\frac{\log p}{p}\left(\log\frac{x}{p^{2}}-2\right)$
$\quad+\sum_{k\geq 3}\sum_{p\leq x^{1/k}}\frac{\log p}{p^{k/2}}\left(\log\frac{x}{p^{k}}-2\right)$
and (2.1), it can be obtained by an argument similar to that of Theorem 4. $\Box$
5. Proof of Theorems 6 and 7
5.1. Proof of Theorem 6
Noting the orthogonality of characters
$\sum_{\chi\,{\rm mod}\,q}\chi(n)=\begin{cases}~\varphi(q)&n\equiv 1\,{\rm mod}\,q,\\ ~0&n\not\equiv 1\,{\rm mod}\,q,\end{cases}$
we take the sum of (4.4) over all $\chi$ modulo $q$. Then,
$\sum_{\chi\,{\rm mod}\,q}f_{\chi}(x)=\varphi(q)\sum_{{n\leq x}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}.$ (5.1)
The Dirichlet theorem on arithmetic progressions [15, Corollary 11.17] tells us that the leading term in the asymptotic formulas for both (5.1) and
$2\,\sum_{\chi\,{\rm mod}\,q}\sum_{n\leq x}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}=2\,\varphi(q)\sum_{{n\leq x}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}$
as $x\to\infty$ is $4\sqrt{x}$. Thus, leading terms cancel out in
$f_{q}(x)$ $=\frac{1}{\varphi(q)}\sum_{\chi\,{\rm mod}\,q}f_{\chi}(x)-\frac{4\sqrt{x}}{\varphi(q)}$ (5.2)
$=\sum_{{n\leq x}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}-\frac{4\sqrt{x}}{\varphi(q)}$
and
$F_{q}(x)$ $=\frac{1}{\varphi(q)}\sum_{\chi\,{\rm mod}\,q}f_{\chi}(x)-\frac{2}{\varphi(q)}\sum_{\chi\,{\rm mod}\,q}\sum_{n\leq x}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}$ (5.3)
$=\frac{1}{\varphi(q)}\sum_{\chi\,{\rm mod}\,q}\sum_{n\leq x}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}\log\frac{x}{ne^{2}}$
$=\sum_{{n\leq x}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{ne^{2}}.$
Taking the sum $\varphi(q)^{-1}\sum_{\chi\,{\rm mod}\,q}$ in (4.5),
$\int_{1}^{\infty}$ $(-f_{q}(x))x^{-s+1/2}\,\frac{dx}{x}=\frac{1}{\varphi(q)}\left(\frac{1}{(s-1/2)^{2}}\sum_{\chi\,{\rm mod}\,q}\frac{L^{\prime}}{L}(s,\chi)+\frac{4}{s-1}\right)$ (5.4)
for $\Re(s)>1$. On the other hand, by (3.6), (4.5), and the way of their proofs, we obtain
$\int_{1}^{\infty}\left(-\sum_{n\leq x}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}\log\frac{x}{ne^{2}}\right)x^{-s+1/2}\,\frac{dx}{x}=\frac{2(1-s)}{(s-1/2)^{2}}\frac{L^{\prime}}{L}(s,\chi)$ (5.5)
for $\Re(s)>1$. Taking the sum $\varphi(q)^{-1}\sum_{\chi\,{\rm mod}\,q}$ in (5.5),
$\int_{1}^{\infty}(-F_{q}(x))x^{-s+1/2}\,\frac{dx}{x}=\frac{2(1-s)}{(s-1/2)^{2}}\frac{1}{\varphi(q)}\sum_{\chi\,{\rm mod}\,q}\frac{L^{\prime}}{L}(s,\chi).$ (5.6)
Then, following an argument similar to the proof of Theorems 2 and 8, and using (5.4) and (5.6), we obtain the first claim of Theorem 6.
We now prove the second and third claims of Theorem 6. Suppose that $L(1/2,\chi)\not=0$ for any Dirichlet character $\chi$ modulo $q$. In an argument similar to the proof of (3.8), in the case $\chi=\chi_{0}$, we include the zeros of $\prod_{p|q}(1-p^{-s})$ in both sums for $\rho$ on the right-hand side of (3.8); in the case $\chi\neq\chi_{0}$, we use (4.8) and (2.13) instead of (3.1) and (2.6), respectively, and include the zeros of $\prod_{p|q}(1-\chi^{\ast}(p)p^{-s})$ in both sums for $\rho$ on the right-hand side, to obtain
$\sum_{n\leq x}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}\log\frac{x}{ne^{2}}$ $=-\frac{L^{\prime}}{L}\left(\frac{1}{2},\chi\right)\log x+2\sum_{\rho_{\chi}}\frac{x^{\rho_{\chi}-1/2}}{\rho_{\chi}-1/2}$ (5.7)
$\quad-\sum_{\rho_{\chi}}\frac{x^{\rho_{\chi}-1/2}}{(\rho_{\chi}-1/2)^{2}}-\left(\frac{L^{\prime}}{L}\right)^{\prime}\left(\frac{1}{2},\chi\right)+2\frac{L^{\prime}}{L}\left(\frac{1}{2},\chi\right)$
$\quad-\sum_{k=0}^{\infty}\frac{x^{-2k-\kappa^{\ast}-1/2}}{(2k+\kappa^{\ast}+1/2)^{2}}-2\sum_{k=0}^{\infty}\frac{x^{-2k-\kappa^{\ast}-1/2}}{2k+\kappa^{\ast}+1/2}$
$\quad+O\left(\frac{\log x}{\sqrt{x}}\right).$
Here, the two sums $\sum_{\rho_{\chi}}$ on the right-hand side are taken over all nontrivial zeros of $L(s,\chi^{\ast})$ and all zeros of the finite product $L(s,\chi)/L(s,\chi^{\ast})=\prod_{p\mid q}(1-\chi^{\ast}(p)p^{-s})$, $\kappa^{\ast}=\kappa(\chi^{\ast})$, and the last term arises from the difference between $\sideset{}{{}^{\prime}}{\sum}_{n\leq x}$ and $\sum_{n\leq x}$. Taking the sum $\sum_{\chi\,{\rm mod}\,q}$ in (4.8) and using (3.1) for $\chi=\chi_{0}$, we obtain
$\varphi(q)f_{q}(x)$ $=-\log x\sum_{\chi\,{\rm mod}\,q}\frac{L^{\prime}}{L}\left(\frac{1}{2},\chi\right)-\sum_{\chi\,{\rm mod}\,q}\sum_{\rho_{\chi^{\ast}}}\frac{x^{\rho_{\chi^{\ast}}-1/2}}{(\rho_{\chi^{\ast}}-1/2)^{2}}+O(1),$ (5.8)
where the sums $\sum_{\rho_{\chi^{\ast}}}$ on the right-hand side are taken over all nontrivial zeros of $L(s,\chi^{\ast})$ and the implied constant depends only on $q$. Similarly, taking the sum $\sum_{\chi\,{\rm mod}\,q}$ in (5.7), we obtain
$\varphi(q)F_{q}(x)$ $=-\log x\sum_{\chi\,{\rm mod}\,q}\frac{L^{\prime}}{L}\left(\frac{1}{2},\chi\right)$ (5.9)
$\quad+\sum_{\chi\,{\rm mod}\,q}\left[2\sum_{\rho_{\chi^{\ast}}}\frac{x^{\rho_{\chi^{\ast}}-1/2}}{\rho_{\chi^{\ast}}-1/2}-\sum_{\rho_{\chi^{\ast}}}\frac{x^{\rho_{\chi^{\ast}}-1/2}}{(\rho_{\chi^{\ast}}-1/2)^{2}}\right]+O(1),$
where the sums $\sum_{\rho_{\chi}^{\ast}}$ and the implied constant have the same meaning as in (5.8). In (5.9), the contribution from the zeros of $L(s,\chi)/L(s,\chi^{\ast})$ is included in the $O(1)$ term, and it is justified as follows.
For a Dirichlet character $\chi$ modulo $q$, we denote by $\mathcal{Z}_{1}(\chi)$ the set of all nontrivial zeros of $L(s,\chi^{\ast})$ counted with multiplicity and by $\mathcal{Z}_{2}(\chi)$ the set of all zeros of $L(s,\chi)/L(s,\chi^{\ast})$. Then, the second sum on the right-hand side of (2.13) for $s=1/2$ is written as
$\sum_{\rho_{\chi}}\frac{x^{\rho_{\chi}-1/2}}{\rho_{\chi}-1/2}=\lim_{T\to\infty}\left(\sum_{{\rho_{\chi}\in\mathcal{Z}_{1}(\chi)}\atop{|\Im(\rho_{\chi})|\leq T}}\frac{x^{\rho_{\chi}-1/2}}{\rho_{\chi}-1/2}+\sum_{{\rho_{\chi}\in\mathcal{Z}_{2}(\chi)}\atop{|\Im(\rho_{\chi})|\leq T}}\frac{x^{\rho_{\chi}-1/2}}{\rho_{\chi}-1/2}\right).$
The second sum inside the brackets on the right-hand side is $x^{-1/2}(1+o(1))$ with respect to $T$, because the zeros of $p$-factors of $L(s,\chi)/L(s,\chi^{\ast})$ have the form $i(A+Bn)$ ($n\in\mathbb{Z}$) for some real numbers $A$ and $B\not=0$, and the series
$\lim_{N\to\infty}\sum_{|n|\leq N}\frac{x^{iBn}}{i(A+Bn)-1/2}$
is convergent and bounded, as in the case of the Fourier series of the sawtooth wave. Therefore, the contribution from the zeros of $L(s,\chi)/L(s,\chi^{\ast})$ on the right-hand side of (5.7), which originates from (2.13), is bounded. Moreover, the contributions from the zeros of $L(s,\chi)/L(s,\chi^{\ast})$ arising from (3.1) and (4.8) are also bounded, since the sums over zeros in (3.1) and (4.8) converge absolutely.
Assuming that the GRH holds for $L(s,\chi)$ for all $\chi$ modulo $q$, the second sum on the right-hand side of (5.8) is bounded. Additionally, if we assume (1.25), the last sum over $\chi$ mod $q$ on the right-hand side of (5.9) is $o(\log x)$. Hence, (1.23) and (1.24) hold.
Conversely, assuming (1.23) (resp. (1.24)), it follows from (5.2) (resp. (5.3)) that $f_{q}(x)\ll\log x$ (resp. $F_{q}(x)\ll\log x$) as $x\to\infty$. Therefore, the integral on the left-hand side of (5.4) (resp. (5.6)) converges uniformly on any compact subset of the right half-plane $\Re(s)>1/2$, and hence defines an analytic function there. Thus, the GRH for $L(s,\chi)$ for all $\chi$ modulo $q$ follows from an argument similar to that in the proof of Theorem 8. Moreover, assuming (1.24), it follows from (5.3) that
$\varphi(q)F_{q}(x)=-\log x\sum_{\chi\,{\rm mod}\,q}\frac{L^{\prime}}{L}\left(\frac{1}{2},\chi\right)+o(\log x)\quad\text{as }x\to\infty.$
Comparing this with (5.9), we find that the sum over $\chi$ mod $q$ on the right-hand side of (5.9) must be $o(\log x)$. However, the latter part of this sum is absolutely convergent and bounded, and thus (1.25) follows.
Finally, we prove the fourth and fifth claims of Theorem 6. By using (2.14) in place of (2.13) and (4.10) in place of (4.8), and employing the resulting modified versions of (5.8) and (5.9), one can show that both (1.26) and (1.27) are equivalent to the GRH for $L(s,\chi)$ for all Dirichlet characters $\chi$ modulo $q$, by an argument similar to that used in the proofs of the second and third claims. Furthermore, by appropriately modifying (5.9) using (2.14) and (4.10), one can also show that (1.27) implies (1.28) by an argument similar to that of the proof of the third claim. We omit the details, as there is no essential difference. $\Box$
5.2. Proof of Theorem 7
First, we prove (1.29). After multiplying $f_{q}(x)$ in (5.2) by $\varphi(q)$, summing over $3\leq q\leq Q$, and then changing the order of summation, we obtain
$\sum_{3\leq q\leq Q}\varphi(q)f_{q}(x)$ $=\sum_{3\leq q\leq Q}\left(\varphi(q)\sum_{{n\leq x}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}-4\sqrt{x}\right)$ (5.10)
$=\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}\sum_{{3\leq q\leq Q}\atop{q\mid n-1}}\varphi(q)-4\sqrt{x}\,(Q-2),$
where we understand that the inner sum is zero if it is an empty sum. For any $n\leq x$, all the divisors of n1n-1 appear in the range $1\leq q\leq Q$ by the assumption $Q\geq x$. For a prime number $p$, 22 divides $p^{k}-1$ if and only if $p$ is odd. Therefore, using $\sum_{d|m}\varphi(d)=m$ on the right-hand side of (5.10),
$\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}\sum_{{3\leq q\leq Q}\atop{q\mid n-1}}\varphi(q)$ $=\sum_{{n\leq x}\atop{n\not\in 2^{\mathbb{N}}}}\frac{\Lambda(n)(n-3)}{\sqrt{n}}\log\frac{x}{n}+\log 2\sum_{2^{k}\leq x}\frac{(2^{k}-2)}{\sqrt{2^{k}}}\log\frac{x}{2^{k}}$
$=\sum_{n\leq x}\frac{\Lambda(n)(n-3)}{\sqrt{n}}\log\frac{x}{n}+o(\sqrt{x}\,).$
Using [21, (2.3)] and the prime number theorem (2.2) in (3.1),
$\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}=O(\sqrt{x}).$
Therefore, we obtain
$\sum_{3\leq q\leq Q}\varphi(q)f_{q}(x)=\sum_{n\leq x}\Lambda(n)\sqrt{n}\,\log\frac{x}{n}-4\sqrt{x}\,Q+O(\sqrt{x}\,).$
By applying partial summation to the first term on the right-hand side using the prime number theorem (2.2),
$\sum_{n\leq x}\Lambda(n)\cdot\sqrt{n}\,\log\frac{x}{n}$ $=-\int_{1}^{x}\sum_{n\leq u}\Lambda(n)\left(-\frac{1}{\sqrt{u}}+\frac{1}{2\sqrt{u}}\log\frac{x}{u}\right)\,du$
$=\frac{4}{9}x\sqrt{x}\,(1+o(1)).$
Hence, we obtain (1.29).
Next, we prove (1.30). After multiplying $F_{q}(x)$ in (5.3) by $\varphi(q)$, summing over $3\leq q\leq Q$, and then changing the order of summation as in (5.10), we obtain
$\sum_{3\leq q\leq Q}\varphi(q)F_{q}(x)=\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{ne^{2}}\sum_{{3\leq q\leq Q}\atop{q\mid n-1}}\varphi(q).$ (5.11)
Then, as before,
$\sum_{3\leq q\leq Q}\varphi(q)F_{q}(x)=\sum_{n\leq x}\frac{\Lambda(n)(n-3)}{\sqrt{n}}\log\frac{x}{ne^{2}}+o(\sqrt{x}\,).$
Using [21, (2.3)] and the prime number theorem (2.2) in (3.8),
$\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{ne^{2}}=o(\sqrt{x}\,).$
Therefore, we obtain
$\sum_{3\leq q\leq Q}\varphi(q)F_{q}(x)=\sum_{n\leq x}\Lambda(n)\sqrt{n}\,\log\frac{x}{ne^{2}}+o(\sqrt{x}\,).$
By applying partial summation to the first term on the right-hand side using the prime number theorem (2.2),
$\sum_{n\leq x}\Lambda(n)\cdot\sqrt{n}\,\log\frac{x}{ne^{2}}$ $=-2\sqrt{x}\,\sum_{n\leq x}\Lambda(n)-\int_{1}^{x}\sum_{n\leq u}\Lambda(n)\left(-\frac{2}{\sqrt{u}}+\frac{1}{2\sqrt{u}}\log\frac{x}{u}\right)\,du$
$=-\frac{8}{9}x\sqrt{x}\,(1+o(1)).$
Hence, we obtain (1.30) by substituting $xe^{2}$ for $x$. $\Box$
If we take the sum $\sum_{3\leq q\leq Q}$ in (5.3), we obtain
$\sum_{3\leq q\leq Q}F_{q}(x)=\sum_{n\leq x}\frac{\Lambda(n)(\sigma_{0}(n-1)-\nu(n))}{\sqrt{n}}\log\frac{x}{ne^{2}},$
where $\sigma_{0}(n)$ denotes the number of divisors of $n$ and $\nu(n)=1$ if $n=2^{k}$ for some $k\in\mathbb{Z}_{>0}$, and $\nu(n)=2$ otherwise. In this case, the study of Titchmarsh’s divisor problem $\sum_{n\leq x}\Lambda(n)\sigma_{0}(n-1)$ provides the leading term. Using the asymptotic formula
$\sum_{n\leq x}\Lambda(n)\sigma_{0}(n-1)$ $=\frac{\zeta(2)\zeta(3)}{\zeta(6)}x\log x$
$\quad+\frac{\zeta(2)\zeta(3)}{\zeta(6)}\left[2\left(C_{0}-\sum_{p}\frac{\log p}{p^{2}-p+1}\right)-1\right]x+O_{A}\left(\frac{x}{(\log x)^{A}}\right)$
obtained by Bombieri, Friedlander, and Iwaniec [4] and Fouvry [6], we find that
$\sum_{n\leq x}\frac{\Lambda(n)\sigma_{0}(n-1)}{\sqrt{n}}\log\frac{x}{ne^{2}}=-8\,\frac{\zeta(2)\zeta(3)}{\zeta(6)}\sqrt{x}\,(1+o(1)).$
This also shows that (1.22) has a constant sign on average, however, the sum used in the proof of Theorem 7 is probably easier to handle.

### v3

@@
 According to the prime number theorem, the leading term in the asymptotic behavior of $\pi(x)$ as $x\to\infty$ is the logarithmic integral ${\rm li}(x)=\int_{0}^{x}dt/\log t$.
-The difference $\pi(x)-{\rm li}(x)$ had long been conjectured to be negative for all $x$.
+The difference $\pi(x)-{\rm li}(x)$ had long been conjectured to be negative for all $x>2$.
 This conjecture was disproven in 1914 by Littlewood, who showed that it changes sign infinitely often [8].
@@
 Furthermore, as mentioned in Pintz [16] and as proved by Johnston [10], the RH is equivalent to the negativity
-$\int_{2}^{x}(\pi(t)-{\rm li}(t))\,dt<0\quad(x>2).$ (1.1)
+$\int_{2}^{x}(\pi(t)-{\rm li}(t))\,dt<0\quad(x>2).$ (1)
 Johnston also shows that the RH is equivalent to the negativity
-$\int_{2}^{x}(\vartheta(t)-t)\,dt<0\quad(x>2),$ (1.2)
+$\int_{2}^{x}(\vartheta(t)-t)\,dt<0\quad(x>2),$ (2)
 where $\vartheta(x)=\sum_{p\leq x}\log p$, the first Chebyshev function.
@@
 On the other hand, a negative bias is observed in the difference between the weighted Chebyshev function
-$\psi_{1/2}(x):=\sideset{}{{}^{\prime}}{\sum}_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}$ (1.3)
+$\psi_{1/2}(x):=\sideset{}{{}^{\prime}}{\sum}_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}$ (3)
 and its leading term $2\sqrt{x}$ as $x\to\infty$. Regarding this observation, using the explicit formula
-$\psi_{1/2}(x)-2\sqrt{x}=-\sum_{\rho}\frac{x^{\rho-1/2}}{\rho-1/2}-\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)+o(1)$ (1.4)
-in (2.6) below, together with Littlewood’s result [8, Theorem 5.6]
+$\psi_{1/2}(x)-2\sqrt{x}=-\sum_{\rho}\frac{x^{\rho-1/2}}{\rho-1/2}-\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)+o(1)$ (4)
+in (36) below, together with Littlewood’s result [8, Theorem 5.6]
 $\sum_{\rho}\frac{x^{\rho}}{\rho}=\Omega_{\pm}(\sqrt{x}\log\log\log x),$
@@
 $\liminf_{x\to\infty}f(x)/g(x)<0$), and $f(x)=\Omega_{\pm}(g(x))$ means that both $f(x)=\Omega_{+}(g(x))$ and $f(x)=\Omega_{-}(g(x))$ hold.
-By the second functional equation in (2.4), if $\rho$ is a nontrivial zero of $\zeta(s)$, then so is $\bar{\rho}$.
-Therefore, both $\sum_{\rho}x^{\rho-1/2}/\rho$ and $\sum_{\rho}x^{\rho-1/2}/(\rho-1/2)$ are real valued.
+By the second functional equation in (34), if $\rho$ is a nontrivial zero of $\zeta(s)$, then so is $\bar{\rho}$.
+Therefore, both $\sum_{\rho}x^{\rho}/\rho$ and $\sum_{\rho}x^{\rho-1/2}/(\rho-1/2)$ are real valued.
 However, assuming the RH, $\psi_{1/2}(x)-2\sqrt{x}$ is negative on average, because
 $\int_{0}^{x}(\psi_{1/2}(y)-2\sqrt{y}\,)\frac{dy}{y}=-\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)\log x-\sum_{\rho}\frac{x^{\rho-1/2}}{(\rho-1/2)^{2}}+o(1)$
-holds, $(\zeta^{\prime}/\zeta)(1/2)=2.68609\dots$ (see (3.1) and (3.2) below), and the sum on the right-hand side is bounded under the RH by $\sum_{\rho}|\rho|^{-2}<\infty$ (see §2.2).
+holds, $(\zeta^{\prime}/\zeta)(1/2)=2.68609\dots$ (see (46) and (47) below), and the sum on the right-hand side is bounded under the RH by $\sum_{\rho}|\rho|^{-2}<\infty$ (see §2.2).
 This bias in the difference $\psi_{1/2}(x)-2\sqrt{x}$ is also a sufficient condition for the RH as in the case of $\pi(x)-{\rm li}(x)$ and $\vartheta(x)-x$:
@@
     There exists an $x_{0}\geq 2$ such that
-    $\int_{0}^{x}$ $(\psi_{1/2}(y)-2\sqrt{y}\,)\frac{dy}{y}$ (1.5)
+    $\int_{0}^{x}$ $(\psi_{1/2}(y)-2\sqrt{y}\,)\frac{dy}{y}$ (5)
     $=\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}-4\sqrt{x}~\leq~0$
@@
     It holds that
-    $\lim_{x\to\infty}\left[\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)-\frac{4\sqrt{x}}{\log x}\right]=-\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right).$ (1.6)
-The prime on the sum is unnecessary on the left-hand side of (1.5), since $\log(x/n)$ vanishes when $n=x$. Taking the average
+    $\lim_{x\to\infty}\left[\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)-\frac{4\sqrt{x}}{\log x}\right]=-\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right).$ (6)
+The prime on the sum is unnecessary on the left-hand side of (5), since $\log(x/n)$ vanishes when $n=x$. Taking the average
 $\frac{1}{x}\int_{1}^{x}\frac{1}{2}\sqrt{\frac{x}{u}}\log\sqrt{\frac{x}{u}}\,du=1-\frac{1}{2\sqrt{x}}(\log x+2)$
-into account, the inequality in (1.5) can be rewritten as an inequality for the sum of $\Lambda(n)$ with a weight that is approximately one on average:
+into account, the inequality in (5) can be rewritten as an inequality for the sum of $\Lambda(n)$ with a weight that is approximately one on average:
 $\sum_{n\leq x}\Lambda(n)\cdot\frac{1}{2}\sqrt{\frac{x}{n}}\log\sqrt{\frac{x}{n}}-x~\leq~0.$
@@
 Considering [19], we are interested in the logarithmic density of $x$ for which $\psi_{1/2}(x)-2\sqrt{x}\leq 0$ holds, but we do not study that topic in this paper.
-The equivalence between the first and second statements in Theorem 1 is quite similar to that between the RH and either (1.1) or (1.2).
-However, an analog of (1.6) for $\pi(x)-\mathrm{li}(x)$ or $\vartheta(x)-x$ does not appear to have been previously noted in the literature.
-Moreover, it should be mentioned that Theorem 1 was not discovered as an analog of (1.1); rather, it arose in connection with the screw functions discussed below.
+The equivalence between the first and second statements in Theorem 1 is quite similar to that between the RH and either (1) or (2).
+However, an analog of (6) for $\pi(x)-\mathrm{li}(x)$ or $\vartheta(x)-x$ does not appear to have been previously noted in the literature.
+Moreover, it should be mentioned that Theorem 1 was not discovered as an analog of (1); rather, it arose in connection with the screw functions discussed below.
 Theorem 1 can be interpreted in terms of the screw function $g_{\zeta}(t)$ for $\zeta(s)$ studied in [21]. It is a function on $\mathbb{R}$ defined by
@@
 with
-$g_{0}(t)$ $=\sum_{n\leq\exp(|t|)}\frac{\Lambda(n)}{\sqrt{n}}(|t|-\log n)-4(e^{t/2}+e^{-t/2}-2),$ (1.7)
+$g_{0}(t)$ $=\sum_{n\leq\exp(|t|)}\frac{\Lambda(n)}{\sqrt{n}}(|t|-\log n)-4(e^{t/2}+e^{-t/2}-2),$ (7)
 $g_{\infty}(t)$ $=-\frac{|t|}{2}\left(\frac{\Gamma^{\prime}}{\Gamma}\left(\frac{1}{4}\right)-\log\pi\right)-\frac{1}{4}\Bigl(\Phi(1,2,1/4)-e^{-|t|/2}\Phi(e^{-2|t|},2,1/4)\Bigr),$
@@
 in [21, Section 2.1], the function $g_{0}(t)$ corresponds to the non-archimedean factor $s(s-1)\zeta(s)$ and $g_{\infty}(t)$ corresponds to the archimedean factor $\pi^{-s/2}\Gamma(s/2)$ of the Riemann xi-function
-$\xi(s):=\frac{1}{2}\,s(s-1)\pi^{-s/2}\Gamma(s/2)\zeta(s).$ (1.8)
+$\xi(s):=\frac{1}{2}\,s(s-1)\pi^{-s/2}\Gamma(s/2)\zeta(s).$ (8)
 According to the Kreĭn–Langer criterion [12, Satz 5.9], $g_{0}(t)$ cannot be a screw function even assuming the RH, because the logarithmic derivative of $s(s-1)\zeta(s)$ can never belong to the Nevanlinna class.
 However, if we note the relation
-$g_{0}(\log x)=\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}-4\left(\sqrt{x}+\frac{1}{\sqrt{x}}-2\right),$ (1.9)
+$g_{0}(\log x)=\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}-4\left(\sqrt{x}+\frac{1}{\sqrt{x}}-2\right),$ (9)
 then we find that the RH is equivalent to $g_{0}(t)$ having a constant sign.
 Corollary 1.
-Let $g_{0}(t)$ be the non-archimedean part of $g_{\zeta}(t)$ defined in (1.7). Then, the RH holds if and only if there exists $t_{0}>0$ such that $g_{0}(t)\geq 0$ holds for all $t\geq t_{0}$.
-Noting (1.4), if we replace $4\sqrt{x}$ in (1.5) with $2\sum_{n\leq x}\Lambda(n)/\sqrt{n}$ and express $2$ as $\log\exp(2)$, we obtain the following as an analog of Theorem 1.
+Let $g_{0}(t)$ be the non-archimedean part of $g_{\zeta}(t)$ defined in (7). Then, the RH holds if and only if there exists $t_{0}>0$ such that $g_{0}(t)\geq 0$ holds for all $t\geq t_{0}$.
+Noting (4), if we replace $4\sqrt{x}$ in (5) with $2\sum_{n\leq x}\Lambda(n)/\sqrt{n}$ and express $2$ as $\log\exp(2)$, we obtain the following as an analog of Theorem 1.
 Theorem 2.
 Suppose that there exists an $x_{0}\geq 2$ such that
-$\sum_{n\leq xe^{2}}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}~\leq~0$ (1.10)
+$\sum_{n\leq xe^{2}}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}~\leq~0$ (10)
 holds for all $x\geq x_{0}$. Then, the RH holds. Furthermore, it holds that
-$\lim_{x\to\infty}\sum_{n\leq xe^{2}}\frac{\Lambda(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)=-\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)$ (1.11)
+$\lim_{x\to\infty}\sum_{n\leq xe^{2}}\frac{\Lambda(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)=-\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)$ (11)
 if and only if the RH holds and the estimate
-$\sum_{\rho}\frac{x^{\rho}}{\rho}=o(\sqrt{x}\log x)$ (1.12)
+$\sum_{\rho}\frac{x^{\rho}}{\rho}=o(\sqrt{x}\log x)$ (12)
 is valid as $x\to\infty$.
-The right-hand side of (1.11) is negative (see (3.2) below). Therefore, (1.11) implies that there exists $x_{0}\geq 2$ such that (1.10) holds for all $x\geq x_{0}$,
-For the sum on the left-hand side of (1.12), the best known estimate under the RH is
+The right-hand side of (11) is negative (see (47) below). Therefore, (11) implies that there exists $x_{0}\geq 2$ such that (10) holds for all $x\geq x_{0}$,
+For the sum on the left-hand side of (12), the best known estimate under the RH is
 $\sum_{\rho}\frac{x^{\rho}}{\rho}\ll\sqrt{x}(\log x)^{2}\quad\text{as $x\to\infty$}$
 [9, pp.
-82–85], but it is not enough to prove (1.10) for large $x$, where the Vinogradov symbol $f(x)\ll g(x)$ mean that there exists a positive constant $M$ such that $|f(x)|\leq Mg(x)$ holds for a prescribed range of $x$.
+82–85], but it is not enough to prove (10) for large $x$, where the Vinogradov symbol $f(x)\ll g(x)$ mean that there exists a positive constant $M$ such that $|f(x)|\leq Mg(x)$ holds for a prescribed range of $x$.
 We use the the Landau symbol $f(x)=O(g(x))$ in the same meaning.
-In the above sense, (1.10) and (1.11) may provide a condition stronger than the RH.
+In the above sense, (10) and (11) may provide a condition stronger than the RH.
 For comparison, we note that Montgomery’s conjecture
 $\limsup_{x\to\infty}\frac{\sum_{\rho}x^{\rho}/\rho}{\sqrt{x}(\log\log\log x)^{2}}=\frac{1}{2\pi}\quad\text{and}\quad\liminf_{x\to\infty}\frac{\sum_{\rho}x^{\rho}/\rho}{\sqrt{x}(\log\log\log x)^{2}}=-\frac{1}{2\pi}$
-[14, p.16] implies the RH and the estimate (1.12).
-The equivalence of (1.11) and (1.12) can be viewed as analogous to Akatsuka [2, Theorem 1], which in turn parallels Conrad [5, Theorem 1.3] for $\zeta(s)$.
+[14, p.16] implies the RH and the estimate (12).
+The equivalence of (11) and (12) can be viewed as analogous to Akatsuka [2, Theorem 1], which in turn parallels Conrad [5, Theorem 1.3] for $\zeta(s)$.
 Taking the logarithm of [2, Corollary 4.4] and using [2, Lemma 2.1], Akatsuka’s result asserts that
 $\lim_{x\to\infty}\left[\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}\log n}-\lim_{\varepsilon\to 0+}\left(\int_{0}^{1-\varepsilon}+\int_{1+\varepsilon}^{\sqrt{x}}\right)\frac{du}{\log u}\right]=\log\left(-\zeta\left(\frac{1}{2}\right)\right)$
-is equivalent to the validity of the RH and (1.12), where $\zeta(1/2)=1.46035\cdots$.
+is equivalent to the validity of the RH and (12), where $\zeta(1/2)=1.46035\cdots$.
 Recently, Akatsuka showed that the equation obtained by replacing the upper limit $\sqrt{x}$ of the integral on the left-hand side with $\sqrt{\vartheta(x)}$ is equivalent to the RH ([3, Theorem 3]).
 A sequence $\{a_{n}\}$ is called Riesz summable of order $k$ to $a$ if
-$\lim_{x\to\infty}\sum_{n\leq x}a_{n}\left(1-\frac{\log n}{\log x}\right)=a$
+$\lim_{x\to\infty}\sum_{n\leq x}a_{n}\left(1-\frac{\log n}{\log x}\right)^{k}=a$
 (cf.
 [15, p.158]).
-From (3.1) below, $\{\Lambda(n)/\sqrt{n}\}$ cannot be Riesz summable of order one even if the RH is assumed.
+From (46) below, $\{\Lambda(n)/\sqrt{n}\}$ cannot be Riesz summable of order one even if the RH is assumed.
 However, Theorem 2 asserts that if we suppose a condition stronger than the RH, then a slightly modified Riesz sum of order one has a limit, and conversely, the existence of such a limit implies the RH.
-If we rewrite (1.10) as
-$\sum_{n\leq x}\Lambda(n)\sqrt{\frac{x}{n}}\log\frac{x}{n}\quad\leq~\sum_{x<n\leq xe^{2}}\Lambda(n)\sqrt{\frac{x}{n}}\log\frac{n}{x}$ (1.13)
+If we rewrite (10) as
+$\sum_{n\leq x}\Lambda(n)\sqrt{\frac{x}{n}}\log\frac{x}{n}\quad\leq~\sum_{x<n\leq xe^{2}}\Lambda(n)\sqrt{\frac{x}{n}}\log\frac{n}{x}$ (13)
 it could be interpreted as a bias depending on the size of the weighted prime powers.
@@
 For any pair of integers $q$ and $a$ with $(q,a)=1$, let $\pi(x;q,a)$ denote the number of primes $p\leq x$ such that $p\equiv a$ mod $q$.
-The leading term in the asymptotic behavior of $\pi(x,q,a)$ as $x\to\infty$ is ${\rm li}(x)/\varphi(q)$, where $\varphi(q)$ is the Euler totient function.
+The leading term in the asymptotic behavior of $\pi(x;q,a)$ as $x\to\infty$ is ${\rm li}(x)/\varphi(q)$, where $\varphi(q)$ is the Euler totient function.
 As in the case of the difference $\pi(x)-{\rm li}(x)$, Chebyshev’s conjecture that $\pi(x;4,1)-\pi(x;4,3)$ is always nonpositive was disproved in [8], but it has been shown that the logarithmic density of the set of $x>2$ for which $\pi(x;4,1)-\pi(x;4,3)$ is negative is close to one [19].
@@
 $\psi_{1/2}(x;q,a):=\sideset{}{{}^{\prime}}{\sum}_{{n\leq x}\atop{n\equiv a\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}$
-for integers $q$ and $a$ with $(q,a)=1$. Then, a negative bias is observed in the difference
+for integers $q$ and $a$ with $(q,a)=1$.
+This is a Chebyshev function version of the weighted counting function $\pi_{1/2}(x;q,a)$ studied in Aoki and Koyama [1].
+Then, a negative bias is observed in the difference
 $\psi_{1/2}(x;4,1)-\psi_{1/2}(x;4,3)=\sideset{}{{}^{\prime}}{\sum}_{n\leq x}\frac{\Lambda(n)\chi_{4}(n)}{\sqrt{n}},$
@@
 Furthermore, we find that $\psi_{1/2}(x;4,1)-\psi_{1/2}(x;4,3)$ is negative on average if and only if all zeros of the Dirichlet $L$-function $L(s,\chi_{4})$ in the critical strip $0\leq\Re(s)\leq 1$ lie on the line $\Re(s)=1/2$.
-The latter is nothing but the Generalized Riemann Hypothesis (GRH) for $L(s,\chi_{4})$ holds.
+The latter is nothing but the Generalized Riemann Hypothesis (GRH) for $L(s,\chi_{4})$.
 Theorem 3.
@@
     There exists an $x_{0}\geq 2$ such that
-    $\int_{0}^{x}(\psi_{1/2}(y;4,1)$ $-\psi_{1/2}(y;4,3))\frac{dy}{y}$ (1.14)
+    $\int_{0}^{x}(\psi_{1/2}(y;4,1)$ $-\psi_{1/2}(y;4,3))\frac{dy}{y}$ (14)
     $=\sum_{n\leq x}\frac{\Lambda(n)\chi_{4}(n)}{\sqrt{n}}\log\frac{x}{n}~\leq~0$
@@
     It holds that
-    $\lim_{x\to\infty}\sum_{n\leq x}\frac{\Lambda(n)\chi_{4}(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)=-\frac{L^{\prime}}{L}\left(\frac{1}{2},\chi_{4}\right).$ (1.15)
-Unlike [5, Theorem 1.3], [2, Theorem 1], and Theorem 2, which include claims stronger than the GRH, it should be noted that (1.15) is equivalent to the GRH for $L(s,\chi_{4})$.
-On the other hand, similar to Theorem 1, the nonpositivity in (1.14) can be interpreted as a property of the non-archimedean part of the screw function associated with $L(s,\chi_{4})$ studied in [22], but we omit the details, as they may be easily inferred.
+    $\lim_{x\to\infty}\sum_{n\leq x}\frac{\Lambda(n)\chi_{4}(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)=-\frac{L^{\prime}}{L}\left(\frac{1}{2},\chi_{4}\right).$ (15)
+Unlike [5, Theorem 1.3], [2, Theorem 1], and Theorem 2, which include claims stronger than the GRH, it should be noted that (15) is equivalent to the GRH for $L(s,\chi_{4})$.
+On the other hand, similar to Theorem 1, the nonpositivity in (14) can be interpreted as a property of the non-archimedean part of the screw function associated with $L(s,\chi_{4})$ studied in [22], but we omit the details, as they may be easily inferred.
 Based on his observation of the negative bias in the difference $\pi(x;4,1)-\pi(x;4,3)$, Chebyshev conjectured
-$\lim_{x\to\infty}\sum_{p>2}(-1)^{(p-1)/2}\cdot e^{-p/x}=-\infty.$ (1.16)
-The sum on the left-hand side of (1.16) can be written as
+$\lim_{x\to\infty}\sum_{p>2}(-1)^{(p-1)/2}\cdot e^{-p/x}=-\infty.$ (16)
+The sum on the left-hand side of (16) can be written as
 $\sum_{p\equiv 1\,{\rm mod}\,4}e^{-p/x}\quad-\sum_{p\equiv 3\,{\rm mod}\,4}e^{-p/x}.$
 Here $e^{-p/x}\to 1$ as $x\to\infty$ termwise.
-Therefore, Chebyshev’s conjecture (1.16) suggests that there are “more” primes of the form $p=4n+3$ than those of the form $p=4n+1$.
-Subsequently, Hardy and Littlewood [8] and Landau [13] proved that (1.16) and
-$\lim_{x\to\infty}\sum_{p>2}(-1)^{(p-1)/2}\log p\cdot e^{-p/x}=$-$\infty$ (1.17)
+Therefore, Chebyshev’s conjecture (16) suggests that there are “more” primes of the form $p=4n+3$ than those of the form $p=4n+1$.
+Subsequently, Hardy and Littlewood [8] and Landau [13] proved that (16) and
+$\lim_{x\to\infty}\sum_{p>2}(-1)^{(p-1)/2}\log p\cdot e^{-p/x}=$-$\infty$ (17)
 are equivalent to the GRH for $L(s,\chi_{4})$.
-Knapowski and Turán [11] showed that replacing the weight function $\exp(-p/x)$ in (1.16) and (1.17) with $\exp(-(\log(p/x))^{2})$ is likewise equivalent to the GRH for $L(s,\chi_{4})$.
+Knapowski and Turán [11] showed that replacing the weight function $\exp(-p/x)$ in (16) and (17) with $\exp(-(\log(p/x))^{2})$ is likewise equivalent to the GRH for $L(s,\chi_{4})$.
 Fujii [7] proved that when $\exp(-p/x)$ is replaced with $\exp(-(p/x)^{\alpha})$ (resp.
-$2K_{0}(2\sqrt{p/x})$), the result in (1.16) (resp.
-(1.17)) is equivalent to the GRH for $L(s,\chi_{4})$ if $0<\alpha<4.19$, and he conjectured that this would hold for general $\alpha>0$ as well.
-Platt and Trudgian [18] extends Fujii’s result for (1.16) to $0<\alpha<20.4$.
-Aoki and Koyama [1] studies the case where $\exp(-p/x)$ is replaced by $p^{-s}$ ($s\geq 0$) in (1.16).
+$2K_{0}(2\sqrt{p/x})$), the result in (16) (resp.
+(17)) is equivalent to the GRH for $L(s,\chi_{4})$ if $0<\alpha<4.19$, and he conjectured that this would hold for general $\alpha>0$ as well.
+Platt and Trudgian [18] extend Fujii’s result for (16) to $0<\alpha<20.4$.
 In addition, a few other weight functions were considered in the works of Hardy–Littlewood [8] and Fujii [7].
-Based on the above developments and Theorem 3, considering the weight function $\sqrt{x/p}\log(x/p)$ in (1.17) leads to the following result.
+Based on the above developments and Theorem 3, considering the weight function $\sqrt{x/p}\log(x/p)\mathbf{1}_{p\leq x}$ in (17) leads to the following result.
 Theorem 4.
 The GRH for $L(s,\chi_{4})$ holds if and only if
-$\lim_{x\to\infty}\sum_{2<p\leq x}(-1)^{(p-1)/2}\log p\cdot\sqrt{\frac{x}{p}}\,\log\frac{x}{p}=-\infty.$ (1.18)
+$\lim_{x\to\infty}\sum_{2<p\leq x}(-1)^{(p-1)/2}\log p\cdot\sqrt{\frac{x}{p}}\,\log\frac{x}{p}=-\infty.$ (18)
 As a result comparable to Theorem 4, we note that the asymptotic formula
-$\sum_{2<p\leq x}(-1)^{(p-1)/2}\sqrt{\frac{x}{p}}~\sim~\frac{1}{2}\sqrt{x}\log\log x\quad\text{as $x\to\infty$}$ (1.19)
-holds if and only if the GRH for $L(s,\chi_{4})$ holds and (1.12) is satisfied for the nontrivial zeros of $L(s,\chi_{4})$ by combining [1, Theorem 1.1, (1.3)] and [5, Theorem 6.2].
-According to the proof of Theorem 4, the GRH for $L(s,\chi_{4})$ is equivalent to the left-hand side of (1.18) being asymptotically equal to $(\sqrt{x}/4)(\log x)^{2}$; thus, unlike (1.19), it does not yield a result stronger than the GRH.
+$\sum_{2<p\leq x}(-1)^{(p-1)/2}\sqrt{\frac{x}{p}}~\sim~\frac{1}{2}\sqrt{x}\log\log x\quad\text{as $x\to\infty$}$ (19)
+holds if and only if the GRH for $L(s,\chi_{4})$ holds and (12) is satisfied for the nontrivial zeros of $L(s,\chi_{4})$ by combining [1, Theorem 1.1, (1.3)] and [5, Theorem 6.2].
+According to the proof of Theorem 4, the GRH for $L(s,\chi_{4})$ is equivalent to the left-hand side of (18) being asymptotically equal to $(\sqrt{x}/4)(\log x)^{2}$; thus, unlike (19), it does not yield a result stronger than the GRH.
 However, an analog of Theorem 4 for $\zeta(s)$ yields a conclusion stronger than the RH, similar to Theorem 2.
@@
 The RH holds assuming
-$\lim_{x\to\infty}\sum_{p\leq xe^{2}}\log p\cdot\sqrt{\frac{x}{p}}\,\log\frac{x}{p}=-\infty.$ (1.20)
-Conversely, if the RH holds and (1.12) is valid as $x\to\infty$, then (1.20) holds.
+$\lim_{x\to\infty}\sum_{p\leq xe^{2}}\log p\cdot\sqrt{\frac{x}{p}}\,\log\frac{x}{p}=-\infty.$ (20)
+Conversely, if the RH holds and (12) is valid as $x\to\infty$, then (20) holds.
 1.3. Results on averages over Dirichlet characters
 As naturally expected, Theorem 3 generalizes to a result for general Dirichlet $L$-functions, as stated in Theorem 8 below.
-This generalization is also described by inequalities for oscillating sums, however, these sums can be rewritten as those with a constant sign in both the first and second halves.
+This generalization is also described by inequalities for oscillating sums.
+By considering the averages of these sums, we obtain the following result.
+As in (13), this average can be rewritten as the difference of two sums, each having a constant sign in the first and second halves.
 For the following results, for an imprimitive Dirichlet character $\chi$, we define the GRH to hold for $L(s,\chi)$ as the GRH to hold for $L(s,\chi^{\ast})$, where $\chi^{\ast}$ is a primitive Dirichlet character inducing $\chi$.
@@
     Suppose that there exists $x_{0}\geq 2$ such that either
-    $\sum_{{n\leq x}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}-\frac{4\sqrt{x}}{\varphi(q)}$ (1.21)
+    $\sum_{{n\leq x}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}-\frac{4\sqrt{x}}{\varphi(q)}$ (21)
     has a constant sign for all $x\geq x_{0}$, or
-    $\sum_{{n\leq xe^{2}}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}$ (1.22)
+    $\sum_{{n\leq xe^{2}}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}$ (22)
     has a constant sign for all $x\geq x_{0}$.
@@
     Suppose that $L(1/2,\chi)\not=0$ for all Dirichlet character $\chi$ modulo $q$. Then, we have
-    $\lim_{x\to\infty}$ $\left(\sum_{{n\leq x}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)-\frac{4\sqrt{x}}{\varphi(q)\log x}\right)$ (1.23)
+    $\lim_{x\to\infty}$ $\left(\sum_{{n\leq x}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)-\frac{4\sqrt{x}}{\varphi(q)\log x}\right)$ (23)
     $\qquad\qquad\qquad\qquad\qquad\quad=-\frac{1}{\varphi(q)}\sum_{\chi\,{\rm mod}\,q}\frac{L^{\prime}}{L}\left(\frac{1}{2},\chi\right)$
@@
     Suppose that $L(1/2,\chi)\not=0$ for all Dirichlet character $\chi$ modulo $q$. Then, we have
-    $\lim_{x\to\infty}\sum_{{n\leq xe^{2}}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)=-\frac{1}{\varphi(q)}\sum_{\chi\,{\rm mod}\,q}\frac{L^{\prime}}{L}\left(\frac{1}{2},\chi\right)$ (1.24)
+    $\lim_{x\to\infty}\sum_{{n\leq xe^{2}}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)=-\frac{1}{\varphi(q)}\sum_{\chi\,{\rm mod}\,q}\frac{L^{\prime}}{L}\left(\frac{1}{2},\chi\right)$ (24)
     if and only if the GRH for $L(s,\chi)$ holds for all Dirichlet characters $\chi$ modulo $q$ and
-    $\sum_{\chi\,{\rm mod}\,q}\sum_{\rho_{\chi^{\ast}}}\frac{x^{\rho_{\chi^{\ast}}}}{\rho_{\chi^{\ast}}}=o(\sqrt{x}\log x)$ (1.25)
+    $\sum_{\chi\,{\rm mod}\,q}\sum_{\rho_{\chi^{\ast}}}\frac{x^{\rho_{\chi^{\ast}}}}{\rho_{\chi^{\ast}}}=o(\sqrt{x}\log x)$ (25)
     holds as $x\to\infty$ for nontrivial zeros of $L(s,\chi^{\ast})$ for all $\chi$ modulo $q$, where $\chi^{\ast}$ is a primitive Dirichlet character that induces $\chi$.
@@
     Let $m_{\chi}$ denote the order of the zero of $L(s,\chi)$ at $s=1/2$. Then, we have
-    $\lim_{x\to\infty}$ $\left(\frac{1}{\log x}\sum_{{n\leq x}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)-\frac{4\sqrt{x}}{\varphi(q)(\log x)^{2}}\right)$ (1.26)
+    $\lim_{x\to\infty}$ $\left(\frac{1}{\log x}\sum_{{n\leq x}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)-\frac{4\sqrt{x}}{\varphi(q)(\log x)^{2}}\right)$ (26)
     $\qquad\qquad\qquad\qquad\qquad\quad=-\frac{1}{2}\sum_{\chi\,{\rm mod}\,q}m_{\chi}$
@@
     Under the notation of the third and fourth statements, we have
-    $\lim_{x\to\infty}\,\frac{1}{\log x}\sum_{{n\leq xe^{2}}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)=-\frac{1}{2}\sum_{\chi\,{\rm mod}\,q}m_{\chi}$ (1.27)
+    $\lim_{x\to\infty}\,\frac{1}{\log x}\sum_{{n\leq xe^{2}}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)=-\frac{1}{2}\sum_{\chi\,{\rm mod}\,q}m_{\chi}$ (27)
     if and only if the GRH for $L(s,\chi)$ holds for all Dirichlet characters $\chi$ modulo $q$ and
-    $\sum_{\chi\,{\rm mod}\,q}\sum_{\rho_{\chi^{\ast}}}\frac{x^{\rho_{\chi^{\ast}}}}{\rho_{\chi^{\ast}}}=o(\sqrt{x}(\log x)^{2})$ (1.28)
+    $\sum_{\chi\,{\rm mod}\,q}\sum_{\rho_{\chi^{\ast}}}\frac{x^{\rho_{\chi^{\ast}}}}{\rho_{\chi^{\ast}}}=o(\sqrt{x}(\log x)^{2})$ (28)
     holds as $x\to\infty$ for nontrivial zeros of $L(s,\chi^{\ast})$ for all $\chi$ modulo $q$.
-If $L(1/2,\chi)\not=0$ for all Dirichlet character $\chi$ modulo $q$, then the right-hand sides of (1.26) and (1.27) vanish.
+If $L(1/2,\chi)\not=0$ for all Dirichlet character $\chi$ modulo $q$, then the right-hand sides of (26) and (27) vanish.
 Hence, the second and fourth claims of Theorem 6, as well as the third and fifth claims, are consistent.
-As for (1.10), nonpositivity of (1.22) can be expressed as an inequality for nonnegative sums
+As for (10), nonpositivity of (22) can be expressed as an inequality for nonnegative sums
 $\sum_{{n\leq x}\atop{n\equiv 1\,{\rm mod}\,q}}\Lambda(n)\sqrt{\frac{x}{n}}\log\frac{x}{n}\quad\leq~\sum_{{x<n\leq xe^{2}}\atop{n\equiv 1\,{\rm mod}\,q}}\Lambda(n)\sqrt{\frac{x}{n}}\log\frac{n}{x}.$
-The previous (1.13) corresponds to the case $q=1$ of this inequality.
-According to the second and third statements of Theorem 6, the sign of the sums (1.21) and (1.22) for sufficiently large $x$, which are assumed in the first statement, should be determined by the constant
+The previous (13) corresponds to the case $q=1$ of this inequality.
+According to the second and third statements of Theorem 6, the sign of the sums (21) and (22) for sufficiently large $x$, which are assumed in the first statement, should be determined by the constant
 $C(q):=\sum_{\chi\,{\rm mod}\,q}\frac{L^{\prime}}{L}\left(\frac{1}{2},\chi\right)\in\mathbb{R}.$
@@
 $C(q)=\frac{q-1}{2}(\log 8\pi+C_{0})-\frac{q-2}{2}\log q+\frac{\log q}{\sqrt{q}-1}$
-from (3.2), (4.11), and (4.12) below, where $C_{0}$ is the Euler–Mascheroni constant. It follows that $C(q)<0$ for all odd primes $\leq 47$, while $C(q)>0$ for all primes $\geq 53$.
-The sign constancy of (1.22) is conditional, however, we find that (1.22) is unconditionally negative on average for all sufficiently large $x>1$.
+from (47), (65), and (66) below, where $C_{0}$ is the Euler–Mascheroni constant. It follows that $C(q)<0$ for all odd primes $\leq 47$, while $C(q)>0$ for all primes $\geq 53$.
+The sign constancy of (21) and (22) is conditional. However, we find that they are unconditionally negative on average for all sufficiently large $x>1$.
 Theorem 7.
 We have
-$\sum_{3\leq q\leq Q}$ $\left(\sum_{\chi\,{\rm mod}\,q}\sum_{n\leq x}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}\log\frac{x}{n}-4\sqrt{x}\right)$ (1.29)
+$\sum_{3\leq q\leq Q}$ $\varphi(q)\left(\sum_{{n\leq x}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}-\frac{4\sqrt{x}}{\varphi(q)}\right)$ (29)
 $\qquad\qquad\qquad\qquad=4\sqrt{x}\,\left[\frac{1}{9}x\,(1+o(1))-Q\right]\quad\text{as $x\to\infty$}$
 and
-$\sum_{3\leq q\leq Q}\sum_{\chi\,{\rm mod}\,q}\sum_{n\leq xe^{2}}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}\log\frac{x}{n}=-\frac{8}{9}\,e^{3}\,x\sqrt{x}\,(1+o(1))\quad\text{as $x\to\infty$}$ (1.30)
+$\sum_{3\leq q\leq Q}\varphi(q)\sum_{{n\leq xe^{2}}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}=-\frac{8}{9}\,e^{3}\,x\sqrt{x}\,(1+o(1))\quad\text{as $x\to\infty$}$ (30)
 if $Q\geq x$. In particular, the left-hand sides are negative for all sufficiently large $x>1$.
-The reason $q=1,2$ are excluded in both (1.29) and (1.30) is that, in those cases, the Dirichlet characters modulo $q$ are only the principal character.
-While it is certainly desirable to obtain the negativity of the left-hand side of (1.29) or (1.30) for a large constant $Q\geq 3$, we are also interested in what nontrivial conclusions can be derived when the negativity of the left-hand side of (1.29) or (1.30) is obtained under a condition weaker than $Q\geq x$ with respect to $Q$.
+The reason $q=1,2$ are excluded in both (29) and (30) is that, in those cases, the Dirichlet characters modulo $q$ are only the principal character.
+While it is certainly desirable to obtain the negativity of the left-hand side of (29) or (30) for a large constant $Q\geq 3$, we are also interested in what nontrivial conclusions can be derived when the negativity of the left-hand side of (29) or (30) is obtained under a condition weaker than $Q\geq x$ with respect to $Q$.
 This remains a subject for future study.
@@
 We use Mertens’ theorem [15, Theorem 2.7] in the form
-$\sum_{p\leq x}\frac{\log p}{p}=\log x+O(1)\quad\text{as $x\to\infty$},$ (2.1)
+$\sum_{p\leq x}\frac{\log p}{p}=\log x+O(1)\quad\text{as $x\to\infty$},$ (31)
 and the classical prime number theorem [15, Theorem 6.9]
-$\psi(x)-x=\sum_{\rho}\frac{x^{\rho}}{\rho}=O(x\exp(-c\sqrt{\log x}))\quad\text{as $x\to\infty$},$ (2.2)
+$\psi(x)-x=\sum_{\rho}\frac{x^{\rho}}{\rho}=O(x\exp(-c\sqrt{\log x}))\quad\text{as $x\to\infty$},$ (32)
 where $\sum_{\rho}=\lim_{T\to\infty}\sum_{|\Im(\rho)|\leq T}$ as before.
@@
 Due to the Euler product formula $\zeta(s)=\prod_{p}(1-p^{-s})^{-1}$ [23, (1.1.2)], the logarithmic derivative of $\zeta(s)$ has the Dirichlet series
-$\frac{\zeta^{\prime}}{\zeta}(s)=\sum_{n=1}^{\infty}\frac{\Lambda(n)}{n^{s}}$ (2.3)
-for $\Re(s)>1$ [23, (1.1.8)]. The Riemann xi-function (1.8) is an entire function satisfying the functional equations
-$\xi(s)=\xi(1-s)\quad\text{and}\quad\xi(s)=\overline{\xi(\bar{s})}$ (2.4)
+$\frac{\zeta^{\prime}}{\zeta}(s)=\sum_{n=1}^{\infty}\frac{\Lambda(n)}{n^{s}}$ (33)
+for $\Re(s)>1$ [23, (1.1.8)]. The Riemann xi-function (8) is an entire function satisfying the functional equations
+$\xi(s)=\xi(1-s)\quad\text{and}\quad\xi(s)=\overline{\xi(\bar{s})}$ (34)
 by [23, (2.6.4)] and $\zeta(\sigma)\in\mathbb{R}$ for $\sigma>1$.
@@
 Because $\xi(s)$ is an entire function of order one [23, Theorem 2.12],
-$\sum_{\rho}|\rho|^{-1-\delta}<\infty\quad\text{if $\delta>0$},$ (2.5)
+$\sum_{\rho}|\rho|^{-1-\delta}<\infty\quad\text{if $\delta>0$},$ (35)
 where $\rho$ runs over all nontrivial zeros of $\zeta(s)$ counted with multiplicity. We often use the explicit formula
-$\sideset{}{{}^{\prime}}{\sum}_{n\leq x}\frac{\Lambda(n)}{n^{s}}$ $=\frac{x^{1-s}}{1-s}-\frac{\zeta^{\prime}}{\zeta}(s)-\sum_{\rho}\frac{x^{\rho-s}}{\rho-s}+\sum_{k=1}^{\infty}\frac{x^{-2k-s}}{2k+s}$ (2.6)
+$\sideset{}{{}^{\prime}}{\sum}_{n\leq x}\frac{\Lambda(n)}{n^{s}}$ $=\frac{x^{1-s}}{1-s}-\frac{\zeta^{\prime}}{\zeta}(s)-\sum_{\rho}\frac{x^{\rho-s}}{\rho-s}+\sum_{k=1}^{\infty}\frac{x^{-2k-s}}{2k+s}$ (36)
 for $x>1$, $s\not=1,\rho,-2k$ ($k\in\mathbb{Z}_{>0}$) [15, Exercises 12.1.1, 4].
@@
 If $\chi$ is not a principal character, then $L(s,\chi)$ extends to an entire function and
-$L(1,\chi)\not=0$ (2.7)
+$L(1,\chi)\not=0$ (37)
 [15, Theorem 4.9]. Due to the Euler product formula
-$L(s,\chi)=\prod_{p}(1-\chi(p)p^{-s})^{-1}\quad$ (2.8)
+$L(s,\chi)=\prod_{p}(1-\chi(p)p^{-s})^{-1}\quad$ (38)
 for $\Re(s)>1$ [15, (4.21)], the logarithmic derivative of $L(s,\chi)$ has the absolutely convergent Dirichlet series
-$\frac{L^{\prime}}{L}(s,\chi)=\sum_{n=1}^{\infty}\frac{\Lambda(n)\chi(n)}{n^{s}}$ (2.9)
+$\frac{L^{\prime}}{L}(s,\chi)=\sum_{n=1}^{\infty}\frac{\Lambda(n)\chi(n)}{n^{s}}$ (39)
 for $\Re(s)>1$ [15, (4.25)].
@@
 and
-$\xi(s,\chi):=\left(\frac{q}{\pi}\right)^{(s+\kappa)/2}\Gamma\left(\frac{s+\kappa}{2}\right)L(s,\chi).$ (2.10)
+$\xi(s,\chi):=\left(\frac{q}{\pi}\right)^{(s+\kappa)/2}\Gamma\left(\frac{s+\kappa}{2}\right)L(s,\chi).$ (40)
 Then, it is an entire function satisfying the functional equation
-$\xi(s,\chi)=\varepsilon(\chi)\xi(1-s,\overline{\chi}),\quad\varepsilon(\chi):=\frac{\tau(\chi)}{i^{\kappa}\sqrt{q}}$ (2.11)
+$\xi(s,\chi)=\varepsilon(\chi)\xi(1-s,\overline{\chi}),\quad\varepsilon(\chi):=\frac{\tau(\chi)}{i^{\kappa}\sqrt{q}}$ (41)
 [15, (10.17) and Corollary 10.8], where $\tau(\chi)$ is the Gauss sum of $\chi$.
@@
 Because $\xi(s,\chi)$ is an entire function of order one by [15, (10.31)],
-$\sum_{\rho_{\chi}}|\rho_{\chi}|^{-1-\delta}<\infty\quad\text{if $\delta>0$},$ (2.12)
-where $\rho_{\chi}$ runs over all nontrivial zeros of $L(s,\chi)$ counted with multiplicity. By an argument similar to the proof of (2.6), we obtain
-$\sideset{}{{}^{\prime}}{\sum}_{n\leq x}\frac{\Lambda(n)\chi(n)}{n^{s}}=-\frac{L^{\prime}}{L}(s,\chi)-\sum_{\rho_{\chi}}\frac{x^{\rho_{\chi}-s}}{\rho_{\chi}-s}+\sum_{k=0}^{\infty}\frac{x^{-2k-\kappa-s}}{2k+\kappa+s}$ (2.13)
+$\sum_{\rho_{\chi}}|\rho_{\chi}|^{-1-\delta}<\infty\quad\text{if $\delta>0$},$ (42)
+where $\rho_{\chi}$ runs over all nontrivial zeros of $L(s,\chi)$ counted with multiplicity. By an argument similar to the proof of (36), we obtain
+$\sideset{}{{}^{\prime}}{\sum}_{n\leq x}\frac{\Lambda(n)\chi(n)}{n^{s}}=-\frac{L^{\prime}}{L}(s,\chi)-\sum_{\rho_{\chi}}\frac{x^{\rho_{\chi}-s}}{\rho_{\chi}-s}+\sum_{k=0}^{\infty}\frac{x^{-2k-\kappa-s}}{2k+\kappa+s}$ (43)
 for $x>1$, $s\not=\rho_{\chi},-2k-\kappa$ ($k\in\mathbb{Z}_{\geq 0}$) [26, (6)].
 Suppose that $L(s,\chi)$ has a zero of order $m\geq 1$ at $s=1/2$.
-By isolating the term of $\rho_{\chi}=1/2$ in the second sum on the right-hand side of (2.13), then taking the limit $s\to 1/2$, we obtain
-$\sideset{}{{}^{\prime}}{\sum}_{n\leq x}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}$ $=-m\log x-\lim_{s\to 1/2}\left[\frac{L^{\prime}}{L}(s,\chi)-\frac{m}{s-1/2}\right]$ (2.14)
+By isolating the term of $\rho_{\chi}=1/2$ in the second sum on the right-hand side of (43), then taking the limit $s\to 1/2$, we obtain
+$\sideset{}{{}^{\prime}}{\sum}_{n\leq x}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}$ $=-m\log x-\lim_{s\to 1/2}\left[\frac{L^{\prime}}{L}(s,\chi)-\frac{m}{s-1/2}\right]$ (44)
 $\quad-\sum_{\rho_{\chi}\not=1/2}\frac{x^{\rho_{\chi}-1/2}}{\rho_{\chi}-1/2}+\sum_{k=0}^{\infty}\frac{x^{-2k-\kappa-1/2}}{2k+\kappa+1/2}.$
@@
 Therefore,
-$L(s,\chi)=L(s,\chi^{\ast})\prod_{p\mid q}(1-\chi^{\ast}(p)p^{-s})$ (2.15)
+$L(s,\chi)=L(s,\chi^{\ast})\prod_{p\mid q}(1-\chi^{\ast}(p)p^{-s})$ (45)
 [15, (9.1), (9.2)].
@@
 3.1. Proof of Theorem 1
-The equality in (1.5) is obtained by integrating (1.3). First, we prove (1.5) and (1.6) assuming the RH. We set
+The equality in (5) is obtained by integrating (3). First, we prove (5) and (6) assuming the RH. We set
 $f(x):=\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}.$
 By applying [20, Lemma 1] to $s=1/2$, we obtain
-$f(x)-4\sqrt{x}$ $=-\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)\log x-\left(\frac{\zeta^{\prime}}{\zeta}\right)^{\prime}\left(\frac{1}{2}\right)$ (3.1)
+$f(x)-4\sqrt{x}$ $=-\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)\log x-\left(\frac{\zeta^{\prime}}{\zeta}\right)^{\prime}\left(\frac{1}{2}\right)$ (46)
 $\quad-\sum_{\rho}\frac{x^{\rho-1/2}}{(\rho-1/2)^{2}}-\sum_{k=1}^{\infty}\frac{x^{-2k-1/2}}{(2k+1/2)^{2}}$
@@
 The second sum on the second line of the right-hand side is clearly bounded for $x\geq 2$.
-The first sum on the second line of the right-hand side is also bounded for $x\geq 2$, because $\rho-1/2$ is purely imaginary by the RH and (2.5).
+The first sum on the second line of the right-hand side is also bounded for $x\geq 2$, because $\rho-1/2$ is purely imaginary by the RH and (35).
 On the other hand,
 $\frac{\xi^{\prime}}{\xi}\left(\frac{1}{2}\right)=\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)+\frac{1}{2}\left[\frac{\Gamma^{\prime}}{\Gamma}\left(\frac{1}{4}\right)-\log\pi\right]=0$
-by the taking logarithmic derivative of the first equation in (2.4) and using $(\Gamma^{\prime}/\Gamma)(1/4)=-\pi/2-3\log 2-C_{0}=-4.22745\dots<0$ ([15, p. 411]). Therefore,
-$\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)=\frac{1}{2}\left(\frac{\pi}{2}+\log 8\pi+C_{0}\right)=2.68609\dots>0.$ (3.2)
-Hence, (1.5) holds for all sufficiently large $x\geq 2$. Furthermore, dividing both sides of (3.1) by $\log x$ and taking the limit as $x\to\infty$, we obtain (1.6).
-By (3.2), equality (1.6) implies that (1.5) holds for all sufficiently large $x$.
-Therefore, it remains to show that the RH holds if there exist $x_{0}\geq 2$ such that (1.5) is valid for all $x\geq x_{0}$.
+by taking the logarithmic derivative of the first equation in (34) and using $(\Gamma^{\prime}/\Gamma)(1/4)=-\pi/2-3\log 2-C_{0}=-4.22745\dots<0$ ([15, p. 411]). Therefore,
+$\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)=\frac{1}{2}\left(\frac{\pi}{2}+\log 8\pi+C_{0}\right)=2.68609\dots>0.$ (47)
+Hence, (5) holds for all sufficiently large $x\geq 2$. Furthermore, dividing both sides of (46) by $\log x$ and taking the limit as $x\to\infty$, we obtain (6).
+By (47), equality (6) implies that (5) holds for all sufficiently large $x$.
+Therefore, it remains to show that the RH holds if there exist $x_{0}\geq 2$ such that (5) is valid for all $x\geq x_{0}$.
 The key point is the integral formula
-$\int_{1}^{\infty}(-f(x))\,x^{-s+1/2}\,\frac{dx}{x}=\frac{1}{(s-1/2)^{2}}\frac{\zeta^{\prime}}{\zeta}(s)\quad\text{for}~\Re(s)>1,$ (3.3)
+$\int_{1}^{\infty}(-f(x))\,x^{-s+1/2}\,\frac{dx}{x}=\frac{1}{(s-1/2)^{2}}\frac{\zeta^{\prime}}{\zeta}(s)\quad\text{for}~\Re(s)>1,$ (48)
 which is derived as follows. For an interval $I\subset\mathbb{R}$, we denote by $\mathbf{1}_{I}$ the indicator function of $I$. Using the integral formula
@@
 if $\Re(s)>1$, where the interchange of the order of integration and summation is justified by the absolute convergence of the series for $\Re(s)>1$ and Fubini’s theorem.
-Therefore, we obtain (3.3) by (2.3).
+Therefore, we obtain (48) by (33).
 On the other hand, $(\zeta^{\prime}/\zeta)(s)$ has no poles in the half-line $(1/2,\infty)$ on the real axis except for the simple pole at $s=1$ with residue 1-1, because $\zeta(s)$ has the simple pole at $s=1$ with residue one and has no zeros in the half-line $(1/2,\infty)$ by the series expansion $(1-2^{1-s})\zeta(s)=\sum_{n=1}^{\infty}(-1)^{n-1}n^{-s}>0$ for $s>0$.
-By the integral formula (3.3) and $\int_{1}^{\infty}\sqrt{x}\cdot x^{-s+1/2}dx/x=1/(s-1)$ for $\Re(s)>1$,
-$\int_{1}^{\infty}$ $(-f(x)+4\sqrt{x}\,)x^{-s+1/2}\,\frac{dx}{x}$ (3.4)
+By the integral formula (48) and $\int_{1}^{\infty}\sqrt{x}\cdot x^{-s+1/2}dx/x=1/(s-1)$ for $\Re(s)>1$,
+$\int_{1}^{\infty}$ $(-f(x)+4\sqrt{x}\,)x^{-s+1/2}\,\frac{dx}{x}$ (49)
 $=\frac{1}{(s-1/2)^{2}}\frac{\zeta^{\prime}}{\zeta}(s)+\frac{4}{s-1}\quad\text{for}~\Re(s)>1.$
-The simple pole of $(s-1/2)^{-2}(\zeta^{\prime}/\zeta)(s)$ at $s=1$ is canceled out by the second term of the right-hand side. If we divide the left-hand side of (3.4) into
+The simple pole of $(s-1/2)^{-2}(\zeta^{\prime}/\zeta)(s)$ at $s=1$ is canceled out by the second term of the right-hand side. If we divide the left-hand side of (49) into
 $\left(\int_{x_{0}}^{\infty}+\int_{1}^{x_{0}}\right)(-f(x)+4\sqrt{x}\,)x^{-s+1/2}\,\frac{dx}{x},$
@@
 Therefore, the integral converges uniformly on any compact subset of the right half-plane $\Re(s)>1/2$, and hence defines an analytic function there.
-This implies, by (3.4), that $(\zeta^{\prime}/\zeta)(s)$ has no poles in the right half-plane $\Re(s)>1/2$, which in turn implies that the RH holds.
+This implies, by (49), that $(\zeta^{\prime}/\zeta)(s)$ has no poles in the right half-plane $\Re(s)>1/2$, which in turn implies that the RH holds.
 $\Box$
 3.2. Proof of Corollary 1
-Assuming the RH, it is shown that $g_{0}(t)$ is nonnegative for all sufficiently large $t>0$ by (1.9) and (3.1) as in the proof of Theorem 1.
+Assuming the RH, it is shown that $g_{0}(t)$ is nonnegative for all sufficiently large $t>0$ by (9) and (46) as in the proof of Theorem 1.
 Therefore, we prove the converse.
-Using (1.9), (3.3), and the integral formula
+Using (9), (48), and the integral formula
 $\int_{0}^{\infty}4(e^{t/2}+e^{-t/2}-2)\,e^{-t(s-1/2)}\,dt=\frac{1}{(s-1/2)^{2}}\,\left(\frac{1}{s-1}+\frac{1}{s}\right)\quad\text{for}~\Re(s)>1,$
 we have
-$\int_{0}^{\infty}(-g_{0}(t))\,e^{-t(s-1/2)}\,dt=\frac{1}{(s-1/2)^{2}}\left(\frac{\zeta^{\prime}}{\zeta}(s)+\frac{1}{s-1}+\frac{1}{s}\right)$ (3.5)
+$\int_{0}^{\infty}(-g_{0}(t))\,e^{-t(s-1/2)}\,dt=\frac{1}{(s-1/2)^{2}}\left(\frac{\zeta^{\prime}}{\zeta}(s)+\frac{1}{s-1}+\frac{1}{s}\right)$ (50)
 for $\Re(s)>1$.
 The simple pole of $(\zeta^{\prime}/\zeta)(s)$ at $s=1$ is canceled out by the second term of the right-hand side.
-Hence, if we use (3.5) instead of (3.4), the RH follows from the nonnegativity of $g_{0}(t)$ for all sufficiently large $t>0$ as in the proof of Theorem 1.
+Hence, if we use (50) instead of (49), the RH follows from the nonnegativity of $g_{0}(t)$ for all sufficiently large $t>0$ as in the proof of Theorem 1.
 $\Box$
 3.3. Proof of Theorem 2
-First, we prove that the nonpositivity (1.10) for all sufficiently large $x>0$ implies the RH. Using the integral formula
+First, we prove that the nonpositivity (10) for all sufficiently large $x>0$ implies the RH. Using the integral formula
 $\int_{1}^{\infty}\frac{1}{\sqrt{n}}\cdot\mathbf{1}_{[n,\infty)}(x)\cdot x^{-s+1/2}\,\frac{dx}{x}=\frac{1}{s-1/2}\,n^{-s}\quad\text{for}~\Re(s)>1/2$
@@
 $\int_{1}^{\infty}\left(\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\right)x^{-s+1/2}\,\frac{dx}{x}$ $=-\frac{1}{s-1/2}\frac{\zeta^{\prime}}{\zeta}(s)$
-for $\Re(s)>1$. Combining this with (3.3) and denoting $2$ as $\log\exp(2)$, we obtain
-$\int_{1}^{\infty}\left(-\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{ne^{2}}\right)x^{-s+1/2}\,\frac{dx}{x}=-\frac{2(s-1)}{(s-1/2)^{2}}\frac{\zeta^{\prime}}{\zeta}(s)\quad\text{for}~\Re(s)>1.$ (3.6)
+for $\Re(s)>1$. Combining this with (48) and denoting $2$ as $\log\exp(2)$, we obtain
+$\int_{1}^{\infty}\left(-\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{ne^{2}}\right)x^{-s+1/2}\,\frac{dx}{x}=-\frac{2(s-1)}{(s-1/2)^{2}}\frac{\zeta^{\prime}}{\zeta}(s)\quad\text{for}~\Re(s)>1.$ (51)
 On the right-hand side, as stated in the proof of Theorem 1, $(\zeta^{\prime}/\zeta)(s)$ has no poles in the half-line $(1/2,\infty)$ except for the simple pole at $s=1$, but it is canceled out by s1s-1.
-Hence, it is shown that the nonpositivity (1.10) for sufficiently large $x>0$ implies the RH as in the proof of Theorem 1.
-We prove the second half of the theorem. Applying (2.6) to $s=1/2$,
-$2\sqrt{x}=\psi_{1/2}(x)+\sum_{\rho}\frac{x^{\rho-1/2}}{\rho-1/2}+\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)-\sum_{k=1}^{\infty}\frac{x^{-2k-1/2}}{2k+1/2}$ (3.7)
-for $x>1$. Substituting (3.7) into (3.1),
-$\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}-2\sideset{}{{}^{\prime}}{\sum}_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}$ $=-\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)\log x+2\sum_{\rho}\frac{x^{\rho-1/2}}{\rho-1/2}$ (3.8)
-$\quad-\sum_{\rho}\frac{x^{\rho-1/2}}{(\rho-1/2)^{2}}-\left(\frac{\zeta^{\prime}}{\zeta}\right)^{\prime}\left(\frac{1}{2}\right)+2\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)$
+Hence, it is shown that the nonpositivity (10) for sufficiently large $x>0$ implies the RH as in the proof of Theorem 1.
+We prove the second half of the theorem. Applying (36) to $s=1/2$,
+$2\sqrt{x}=\psi_{1/2}(x)+\sum_{\rho}\frac{x^{\rho-1/2}}{\rho-1/2}+\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)-\sum_{k=1}^{\infty}\frac{x^{-2k-1/2}}{2k+1/2}$ (52)
+for $x>1$. Substituting (52) into (46),
+$\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}-2\sideset{}{{}^{\prime}}{\sum}_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}$ $=-\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)\log x$ (53)
+$\quad+2\sum_{\rho}\frac{x^{\rho-1/2}}{\rho-1/2}-\sum_{\rho}\frac{x^{\rho-1/2}}{(\rho-1/2)^{2}}$
+$\quad-\left(\frac{\zeta^{\prime}}{\zeta}\right)^{\prime}\left(\frac{1}{2}\right)+2\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)$
 $\quad-\sum_{k=1}^{\infty}\frac{x^{-2k-1/2}}{(2k+1/2)^{2}}-2\sum_{k=1}^{\infty}\frac{x^{-2k-1/2}}{2k+1/2}.$
-By denoting $2$ as $\log\exp(2)$, the left-hand side of (3.8) is written as
+By denoting $2$ as $\log\exp(2)$, the left-hand side of (53) is written as
 $\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{ne^{2}}+O\left(\frac{1}{\sqrt{x}}\log x\right),$
 where the error term arises by replacing $\sideset{}{{}^{\prime}}{\sum}_{n\leq x}$ with $\sum_{n\leq x}$.
-Substituting this into (3.8), and then replacing $x$ with $xe^{2}$ and dividing by $\log x$, we have
-$\sum_{n\leq xe^{2}}\frac{\Lambda(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)$ $=-\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)+\frac{2}{\log x}\sum_{\rho}\frac{(xe^{2})^{\rho-1/2}}{\rho-1/2}$ (3.9)
+Substituting this into (53), and then replacing $x$ with $xe^{2}$ and dividing by $\log x$, we have
+$\sum_{n\leq xe^{2}}\frac{\Lambda(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)$ $=-\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)+\frac{2}{\log x}\sum_{\rho}\frac{(xe^{2})^{\rho-1/2}}{\rho-1/2}$ (54)
 $\quad-\frac{1}{\log x}\sum_{\rho}\frac{(xe^{2})^{\rho-1/2}}{(\rho-1/2)^{2}}+O\left(\frac{1}{\log x}\right)$
 as $x\to\infty$.
-Suppose that equation (1.11) holds.
-Then, inequality (1.10) is valid for all sufficiently large $x>0$ by (3.2), and therefore, the RH holds from the first half of the theorem.
-From the RH, the third term on the right hand side of (3.9) is $O(1/\log x)$, and
+Suppose that equation (11) holds.
+Then, inequality (10) is valid for all sufficiently large $x>0$ by (47), and therefore, the RH holds from the first half of the theorem.
+From the RH, the third term on the right hand side of (54) is $O(1/\log x)$, and
 $\sum_{\rho}\frac{x^{\rho-1/2}}{\rho-1/2}-\sum_{\rho}\frac{x^{\rho-1/2}}{\rho}=\frac{1}{2}\sum_{\rho}\frac{x^{\rho-1/2}}{\rho(\rho-1/2)}=O(1).$
-Hence, (1.11) implies that (1.12) holds. Conversely, assuming that the RH and (1.12) hold, (1.11) follows from (3.9). $\Box$
+Hence, (11) implies that (12) holds. Conversely, assuming that the RH and (12) hold, (11) follows from (54). $\Box$
 4. Proof of Theorems 3, 4, and 5
@@
 Let $\chi$ be a non-principal Dirichlet character of modulo $q$, and denote $L(s,\chi)$ by $L(s)$. Suppose that there exists $x_{0}\geq 2$ such that
-$\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\Re(\chi(n))\log\frac{x}{n}$ (4.1)
+$\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\Re(\chi(n))\log\frac{x}{n}$ (55)
 has a constant sign for all $x\geq x_{0}$, and $L(\sigma)\not=0$ for $\sigma>\beta$ for some $1/2\leq\beta<1$.
 Then, $L(s)\not=0$ in the right-half plane $\Re(s)>\beta$.
-In particular, if (4.1) has a constant sign for all sufficiently large $x>0$, and $L(\sigma)\not=0$ for $\sigma>1/2$, then the GRH for $L(s)$ holds.
+In particular, if (55) has a constant sign for all sufficiently large $x>0$, and $L(\sigma)\not=0$ for $\sigma>1/2$, then the GRH for $L(s)$ holds.
 Suppose that $L(1/2)\not=0$. Then, we have
-$\lim_{x\to\infty}\sum_{n\leq x}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)=-\frac{L^{\prime}}{L}\left(\frac{1}{2}\right)$ (4.2)
-if and only if the GRH for $L(s,\chi)$ holds. If $\chi$ is primitive, then equation (4.2) implies that the sum (4.1) has a constant sign for all sufficiently large $x>0$.
+$\lim_{x\to\infty}\sum_{n\leq x}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)=-\frac{L^{\prime}}{L}\left(\frac{1}{2}\right)$ (56)
+if and only if the GRH for $L(s)$ holds. If $\chi$ is primitive, then equation (56) implies that the sum (55) has a constant sign for all sufficiently large $x>0$.
 Furthermore, suppose that $s=1/2$ is a zero of $L(s)$ of order $m\geq 0$. Then, we have
-$\lim_{x\to\infty}\frac{1}{\log x}\sum_{n\leq x}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)=-\frac{m}{2}$ (4.3)
-if and only if the GRH for $L(s,\chi)$ holds. Equation (4.3) implies that the value of the sum (4.1) is negative for all sufficiently large $x>0$ if $m\geq 1$.
+$\lim_{x\to\infty}\frac{1}{\log x}\sum_{n\leq x}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)=-\frac{m}{2}$ (57)
+if and only if the GRH for $L(s)$ holds. Equation (57) implies that the value of the sum (55) is negative for all sufficiently large $x>0$ if $m\geq 1$.
 Remark.
-It is conjectured that $L(\sigma,\chi)\not=0$ for all $1/2\leq\sigma\leq 1$ and for all $\chi$.
-This conjecture is confirmed for primitive Dirichlet characters of modulus $q\leq 4\cdot 10^{5}$ by Platt [17], and for real odd Dirichlet characters of modulus $q\leq 3\cdot 10^{8}$ by Watkins [24].
+It is conjectured that $L(\sigma,\chi)\neq 0$ for all $1/2\leq\sigma\leq 1$ and for all $\chi$.
+This conjecture has been confirmed for primitive Dirichlet characters of modulus $q\leq 4\cdot 10^{5}$ by Platt [17], and for real odd Dirichlet characters of modulus $q\leq 3\cdot 10^{8}$ by Watkins [24].
+It has also been confirmed by Platt [17] that $L(1/2,\chi)\neq 0$ for primitive Dirichlet characters of modulus $q\leq 2\cdot 10^{6}$.
 Proof.
@@
 We set
-$f_{\chi}(x):=\sum_{n\leq x}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}\log\frac{x}{n}.$ (4.4)
+$f_{\chi}(x):=\sum_{n\leq x}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}\log\frac{x}{n}.$ (58)
 Then, the integral formula
-$\int_{1}^{\infty}(-f_{\chi}(x))\,x^{-s+1/2}\,\frac{dx}{x}=\frac{1}{(s-1/2)^{2}}\frac{L^{\prime}}{L}(s)\quad\text{for}~\Re(s)>1$ (4.5)
-is obtained in the same way as (3.3) using the series expansion (2.9).
-On the right-hand side, there exists $1/2\leq\beta<1$ such that $L(\sigma)\not=0$ for any real number $\sigma>\beta$, because $L(s)$ is analytic in $\mathbb{C}$, and has no poles and zeros in the half-line $1,\infty)\subset\mathbb{R}$ by the Euler product (2.8) and Dirichlet’s theorem (2.7).
+$\int_{1}^{\infty}(-f_{\chi}(x))\,x^{-s+1/2}\,\frac{dx}{x}=\frac{1}{(s-1/2)^{2}}\frac{L^{\prime}}{L}(s)\quad\text{for}~\Re(s)>1$ (59)
+is obtained in the same way as (48) using the series expansion (39).
+On the right-hand side, there exists $1/2\leq\beta<1$ such that $L(\sigma)\not=0$ for any real number $\sigma>\beta$, because $L(s)$ is analytic in $\mathbb{C}$, and has no poles and zeros in the half-line $1,\infty)\subset\mathbb{R}$ by the Euler product (38) and Dirichlet’s theorem (37).
 Therefore, if $\chi$ is real, the first half of the theorem follows similarly to the proof of Theorem 1.
-We consider the case where $\chi$ is not real. We replace $\chi$ by $\bar{\chi}$ in (4.5). Combining this with (4.5), we have
-$\int_{1}^{\infty}\left(-\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\Re(\chi(n))\log\frac{x}{n}\right)$ $x^{-s+1/2}\,\frac{dx}{x}$ (4.6)
-$=\frac{1}{(s-1/2)^{2}}\left(\frac{L^{\prime}}{L}(s,\chi)+\frac{L^{\prime}}{L}(s,\bar{\chi})\right).$
+We consider the case where $\chi$ is not real. We replace $\chi$ by $\bar{\chi}$ in (59). Combining this with (59), we have
+$\int_{1}^{\infty}\left(-\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\Re(\chi(n))\log\frac{x}{n}\right)$ $x^{-s+1/2}\,\frac{dx}{x}$ (60)
+$=\frac{1}{2(s-1/2)^{2}}\left(\frac{L^{\prime}}{L}(s,\chi)+\frac{L^{\prime}}{L}(s,\bar{\chi})\right).$
 By $L(s,\bar{\chi})=\overline{L(\bar{s},\chi)}$, the assumption that $L(\sigma,\chi)\neq 0$ for $\sigma>\beta$ implies $L(\sigma,\bar{\chi})\neq 0$ for $\sigma>\beta$.
-This yields that the right-hand side of (4.6) has no poles on the interval $(\beta,\infty)\subset\mathbb{R}$.
-Thus, the assumption on (4.1) implies that the abscissa of convergence for the integral on the left-hand side of (4.6) is at most $\beta$ thanks to Proposition 1.
-This means that the right-hand side of (4.6) is analytic in $\Re(s)>\beta$.
+This yields that the right-hand side of (60) has no poles on the interval $(\beta,\infty)\subset\mathbb{R}$.
+Thus, the assumption on (55) implies that the abscissa of convergence for the integral on the left-hand side of (60) is at most $\beta$ thanks to Proposition 1.
+This means that the right-hand side of (60) is analytic in $\Re(s)>\beta$.
 This forces $L(s,\chi)\neq 0$ in $\Re(s)>\beta$, since the residues of the logarithmic derivative at the zeros are positive integers and hence cannot cancel between $(L^{\prime}/L)(s,\chi)$ and $(L^{\prime}/L)(s,\bar{\chi})$.
 We prove the second half of the theorem. Let $\chi^{\ast}$ be a primitive character that induces $\chi$. Then, by applying the argument in the proof of [20, Lemma 1] to $L(s)$, we obtain
-$\sum_{n\leq x}\frac{\Lambda(n)\chi(n)}{n^{s}}\log\frac{x}{n}$ $=-\frac{L^{\prime}}{L}(s)\log x-\sum_{\rho}\frac{x^{\rho-s}}{(\rho-s)^{2}}$ (4.7)
+$\sum_{n\leq x}\frac{\Lambda(n)\chi(n)}{n^{s}}\log\frac{x}{n}$ $=-\frac{L^{\prime}}{L}(s)\log x-\sum_{\rho}\frac{x^{\rho-s}}{(\rho-s)^{2}}$ (61)
 $\quad-\left(\frac{L^{\prime}}{L}\right)^{\prime}(s)-\sum_{k=0}^{\infty}\frac{x^{-2k-\kappa^{\ast}-s}}{(2k+\kappa^{\ast}+s)^{2}}$
-for $x>1$, $s\not=\rho_{\chi},-2k-\kappa$ ($k\in\mathbb{Z}_{\geq 0}$). If $L(1/2)\not=0$, that is, if $m=0$, then substituting $s=1/2$ into (4.7) yields
-$f_{\chi}(x)$ $=-\frac{L^{\prime}}{L}\left(\frac{1}{2}\right)\log x-\sum_{\rho}\frac{x^{\rho-1/2}}{(\rho-1/2)^{2}}$ (4.8)
+for $x>1$, $s\not=\rho_{\chi},-2k-\kappa^{\ast}$ ($k\in\mathbb{Z}_{\geq 0}$). If $L(1/2)\not=0$, that is, if $m=0$, then substituting $s=1/2$ into (61) yields
+$f_{\chi}(x)$ $=-\frac{L^{\prime}}{L}\left(\frac{1}{2}\right)\log x-\sum_{\rho}\frac{x^{\rho-1/2}}{(\rho-1/2)^{2}}$ (62)
 $\quad-\left(\frac{L^{\prime}}{L}\right)^{\prime}\left(\frac{1}{2}\right)-\sum_{k=0}^{\infty}\frac{x^{-2k-\kappa^{\ast}-1/2}}{(2k+\kappa^{\ast}+1/2)^{2}},$
@@
 Suppose that $L(s)$ has a zero of order $m\geq 1$ at $s=1/2$. Then, we have
-$\lim_{s\to 1/2}$ $\left(-\frac{L^{\prime}}{L}(s)\log x-\frac{x^{-s+1/2}}{(-s+1/2)^{2}}-\left(\frac{L^{\prime}}{L}\right)^{\prime}(s)\right)$ (4.9)
+$\lim_{s\to 1/2}$ $\left(-\frac{L^{\prime}}{L}(s)\log x-\frac{mx^{-s+1/2}}{(-s+1/2)^{2}}-\left(\frac{L^{\prime}}{L}\right)^{\prime}(s)\right)$ (63)
 $=-\frac{m}{2}(\log x)^{2}-\lim_{s\to 1/2}\left[\frac{L^{\prime}}{L}(s,\chi)-\frac{m}{s-1/2}\right]\,\log x$
 $\quad-\lim_{s\to 1/2}\left[\left(\frac{L^{\prime}}{L}\right)^{\prime}(s)+\frac{m}{(s-1/2)^{2}}\right].$
-Therefore, by isolating the term of $\rho=1/2$ in the second sum on the right-hand side of (4.7), then taking the limit $s\to 1/2$, we obtain
-$f_{\chi}(x)$ $=-\frac{m}{2}(\log x)^{2}-\lim_{s\to 1/2}\left[\frac{L^{\prime}}{L}(s,\chi)-\frac{m}{s-1/2}\right]\,\log x$ (4.10)
+Therefore, by isolating the term of $\rho=1/2$ in the second sum on the right-hand side of (61), then taking the limit $s\to 1/2$, we obtain
+$f_{\chi}(x)$ $=-\frac{m}{2}(\log x)^{2}-\lim_{s\to 1/2}\left[\frac{L^{\prime}}{L}(s,\chi)-\frac{m}{s-1/2}\right]\,\log x$ (64)
 $\quad-\lim_{s\to 1/2}\left[\left(\frac{L^{\prime}}{L}\right)^{\prime}(s)+\frac{m}{(s-1/2)^{2}}\right]$
-$\quad-\sum_{\rho\not=1/2}\frac{x^{\rho-1/2}}{(\rho-1/2)^{2}}-\sum_{k=0}^{\infty}\frac{x^{-2k-\kappa^{\ast}-1/2}}{(2k+\kappa^{\ast}+1/2)^{2}},$
-Assuming the GRH for $L(s)$, each term of the right-hand side (4.8) is bounded except for the first term, and each term of the right-hand side (4.10) is bounded except for the first and second terms.
-Hence, (4.2) and (4.3) holds.
-Conversely, assuming (4.2) (resp.
-(4.3)), we have $f_{\chi}(x)\ll\log x$ (resp.
+$\quad-\sum_{\rho\not=1/2}\frac{x^{\rho-1/2}}{(\rho-1/2)^{2}}-\sum_{k=0}^{\infty}\frac{x^{-2k-\kappa^{\ast}-1/2}}{(2k+\kappa^{\ast}+1/2)^{2}}.$
+Assuming the GRH for $L(s)$, each term of the right-hand side (62) is bounded except for the first term, and each term of the right-hand side (64) is bounded except for the first and second terms.
+Hence, (56) and (57) hold.
+Conversely, assuming (56) (resp.
+(57)), we have $f_{\chi}(x)\ll\log x$ (resp.
 $\ll(\log x)^{2}$) as $x\to\infty$.
-Therefore, the integral on the left-hand side of (4.5) converges uniformly on any compact subset in $\Re(s)>1/2$.
+Therefore, the integral on the left-hand side of (59) converges uniformly on any compact subset in $\Re(s)>1/2$.
 Thus, it defines an analytic function in $\Re(s)>1/2$, and hence, $L(s)$ has no zeros (or poles) in $\Re(s)>1/2$, that is the GRH for $L(s)$.
-Finally, we prove that the sum (4.1) has a constant sign for all sufficiently large $x>0$ assuming $L(1/2)\neq 0$, (4.2) and that $\chi$ is primitive.
-Taking the logarithmic derivative of (2.11) and substituting $s=1/2$, we obtain $\Re(\xi^{\prime}/\xi)(1/2,\chi)=0$.
+Finally, we prove that the sum (55) has a constant sign for all sufficiently large $x>0$ assuming $L(1/2)\neq 0$, (56) and that $\chi$ is primitive.
+Taking the logarithmic derivative of (41) and substituting $s=1/2$, we obtain $\Re(\xi^{\prime}/\xi)(1/2,\chi)=0$.
 This gives
-$\Re\frac{L^{\prime}}{L}\left(\frac{1}{2}\right)=\frac{1}{2}\left[\log\frac{\pi}{q}-\frac{\Gamma^{\prime}}{\Gamma}\left(\frac{1}{4}+\frac{\kappa(\chi)}{2}\right)\right]$ (4.11)
-by (2.10). In particular, $\Re(L^{\prime}/L)(1/2)$ depends only on $q$ and the parity of $\chi$. By (4.11), the value $\Re(L^{\prime}/L)(1/2)$ vanishes only when $q>1$ takes the following values
+$\Re\frac{L^{\prime}}{L}\left(\frac{1}{2}\right)=\frac{1}{2}\left[\log\frac{\pi}{q}-\frac{\Gamma^{\prime}}{\Gamma}\left(\frac{1}{4}+\frac{\kappa(\chi)}{2}\right)\right]$ (65)
+by (40). In particular, $\Re(L^{\prime}/L)(1/2)$ depends only on $q$ and the parity of $\chi$. By (65), the value $\Re(L^{\prime}/L)(1/2)$ vanishes only when $q>1$ takes the following values
 $q=\pi\cdot\exp\left[-\frac{\Gamma^{\prime}}{\Gamma}\left(\frac{1}{4}+\frac{\kappa(\chi)}{2}\right)\right]=\begin{cases}~215.332\cdots,&\kappa=0,\\ ~9.305\cdots,&\kappa=1,\end{cases}$
 however it is impossible for any integer $q>1$, where
-$\frac{\Gamma^{\prime}}{\Gamma}\left(\frac{1}{4}+\frac{\kappa(\chi)}{2}\right)=\left(\kappa(\chi)-\frac{1}{2}\right)\pi-3\log 2-C_{0}$ (4.12)
-([15, p. 411]). Therefore, $\Re(L^{\prime}/L)(1/2)$ is a nonzero real number. Hence, by taking the real part of (4.2), it follows that (4.1) has a constant sign for all sufficiently large $x>0$. ∎
+$\frac{\Gamma^{\prime}}{\Gamma}\left(\frac{1}{4}+\frac{\kappa(\chi)}{2}\right)=\left(\kappa(\chi)-\frac{1}{2}\right)\pi-3\log 2-C_{0}$ (66)
+([15, p. 411]). Therefore, $\Re(L^{\prime}/L)(1/2)$ is a nonzero real number. Hence, by taking the real part of (56), it follows that (55) has a constant sign for all sufficiently large $x>0$. ∎
 Proof of Theorem 3.
@@
 Let $\chi$ be a real Dirichlet character such that $L(\sigma,\chi)\not=0$ for $\sigma>\beta~(\geq 1/2)$. Suppose that
-$\sum_{p\leq x}\chi(p)\log p\cdot\sqrt{\frac{x}{p}}\,\log\frac{x}{p}\leq 0$ (4.13)
+$\sum_{p\leq x}\chi(p)\log p\cdot\sqrt{\frac{x}{p}}\,\log\frac{x}{p}\leq 0$ (67)
 for all sufficiently large $x>1$. Then, $L(s,\chi)\not=0$ in the right-half plane $\Re(s)>\beta$. In particular, if $\beta=1/2$, the GRH for $L(s,\chi)$ holds.
 Conversely, suppose that the GRH for $L(s,\chi)$ holds and $L(1/2,\chi)\neq 0$. Then, the asymptotic formula
-$\sum_{p\leq x}\chi(p)\log p\cdot\sqrt{\frac{x}{p}}\,\log\frac{x}{p}=-\frac{\sqrt{x}}{4}(\log x)^{2}\left(1+O\left(\frac{1}{\log x}\right)\right)$ (4.14)
-holds as $x\to\infty$. Therefore, (4.13) is valid for all sufficiently large $x>1$, and the left-hand side of (4.13) tends to -$\infty$ as $x\to\infty$.
+$\sum_{p\leq x}\chi(p)\log p\cdot\sqrt{\frac{x}{p}}\,\log\frac{x}{p}=-\frac{\sqrt{x}}{4}(\log x)^{2}\left(1+O\left(\frac{1}{\log x}\right)\right)$ (68)
+holds as $x\to\infty$. Therefore, (67) is valid for all sufficiently large $x>1$, and the left-hand side of (67) tends to -$\infty$ as $x\to\infty$.
 Proof.
-Let $f_{\chi}(x)$ be the function in (4.4). If we define
+Let $f_{\chi}(x)$ be the function in (58). If we define
 $f_{1}(x):=\sum_{p\leq x}\frac{\chi(p)\log p}{\sqrt{p}}\log\frac{x}{p},$
@@
 $=\frac{1}{2}(\log x)^{2}+O(\log x)-2\sum_{p\leq\sqrt{x}}\frac{(\log p)^{2}}{p}$
-by using (2.1). For the third term on the right-hand side,
+by using (31). For the third term on the right-hand side,
 $\sum_{p\leq x}\frac{(\log p)^{2}}{p}$ $=\log x\sum_{p\leq x}\frac{\log p}{p}-\int_{1}^{x}\left(\sum_{p\leq t}\frac{\log p}{p}\right)\frac{1}{t}\,dt$
@@
 $=\frac{1}{2}(\log x)^{2}+O(\log x)$
-by partial summation and (2.1). Therefore,
-$f_{2}(x)=\frac{1}{4}(\log x)^{2}+O(\log x).$ (4.15)
+by partial summation and (31). Therefore,
+$f_{2}(x)=\frac{1}{4}(\log x)^{2}+O(\log x).$ (69)
 On the other hand,
-$f_{3}(x)=O(\log x)$ (4.16)
-by the convergence of the sum. As a consequence, if we rewrite (4.5) as
-$-\int_{1}^{\infty}f_{1}(x)\,x^{-s+1/2}\frac{dx}{x}$ $=\frac{1}{(s-1/2)^{2}}\frac{L^{\prime}}{L}(s)$ (4.17)
+$f_{3}(x)=O(\log x)$ (70)
+by the convergence of the sum. As a consequence, if we rewrite (59) as
+$-\int_{1}^{\infty}f_{1}(x)\,x^{-s+1/2}\frac{dx}{x}$ $=\frac{1}{(s-1/2)^{2}}\frac{L^{\prime}}{L}(s)$ (71)
 $\quad+\int_{1}^{\infty}f_{2}(x)\,x^{-s+1/2}\frac{dx}{x}+\int_{1}^{\infty}f_{3}(x)\,x^{-s+1/2}\frac{dx}{x},$
@@
 Suppose that $f_{1}(x)$ has a constant sign for all sufficiently large $x>1$.
-Then, the assumption $L(\sigma,\chi)\not=0$ for $\sigma>\beta~(\geq 1/2)$ implies that the abscissa of convergence $\sigma_{c}$ for the integral on the left-hand side of (4.17) satisfies $\sigma_{c}\leq\beta$ by Proposition 1.
-Therefore, the integral on the left-hand side of (4.17) converges uniformly for any compact subset of $\Re(s)>\sigma_{c}$.
-This means that the right-hand side of (4.17) is analytic in $\Re(s)>\beta~(\geq 1/2)$.
+Then, the assumption $L(\sigma,\chi)\not=0$ for $\sigma>\beta~(\geq 1/2)$ implies that the abscissa of convergence $\sigma_{c}$ for the integral on the left-hand side of (71) satisfies $\sigma_{c}\leq\beta$ by Proposition 1.
+Therefore, the integral on the left-hand side of (71) converges uniformly for any compact subset of $\Re(s)>\sigma_{c}$.
+This means that the right-hand side of (71) is analytic in $\Re(s)>\beta~(\geq 1/2)$.
 This forces $L(s,\chi)\neq 0$ in $\Re(s)>\beta~(\geq 1/2)$.
-Conversely, suppose that the GRH for $L(s)$ holds.
-Then, $f_{\chi}(x)=O(\log x)$ by (4.8).
-On the other hand, $f_{\chi}(x)=f_{1}(z)+(1/4)(\log x)^{2}+O(\log x)$ by (4.15) and (4.16).
-Combining these two,
+Conversely, suppose that the GRH for $L(s)$ holds. Then, $f_{\chi}(x)=O(\log x)$ by (62). On the other hand, $f_{\chi}(x)=f_{1}(z)+(1/4)(\log x)^{2}+O(\log x)$ by (69) and (70). Combining these two,
 $f_{1}(x)=-\frac{1}{4}(\log x)^{2}\left(1+O\left(\frac{1}{\log x}\right)\right).$
-This immediately implies (4.14). ∎
+This immediately implies (68). ∎
 4.3. Proof of Theorem 5
-It suffices to show (1.20) with $x$ replaced by $x/e^{2}$. Using the decomposition
+It suffices to show (20) with $x$ replaced by $x/e^{2}$. Using the decomposition
 $\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{ne^{2}}$ $=\sum_{p\leq x}\frac{\log p}{\sqrt{p}}\left(\log\frac{x}{p}-2\right)$
@@
 $\quad+\sum_{k\geq 3}\sum_{p\leq x^{1/k}}\frac{\log p}{p^{k/2}}\left(\log\frac{x}{p^{k}}-2\right)$
-and (2.1), it can be obtained by an argument similar to that of Theorem 4. $\Box$
+and (31), it can be obtained by an argument similar to that of Theorem 4. $\Box$
 5. Proof of Theorems 6 and 7
@@
 $\sum_{\chi\,{\rm mod}\,q}\chi(n)=\begin{cases}~\varphi(q)&n\equiv 1\,{\rm mod}\,q,\\ ~0&n\not\equiv 1\,{\rm mod}\,q,\end{cases}$
-we take the sum of (4.4) over all $\chi$ modulo $q$. Then,
-$\sum_{\chi\,{\rm mod}\,q}f_{\chi}(x)=\varphi(q)\sum_{{n\leq x}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}.$ (5.1)
-The Dirichlet theorem on arithmetic progressions [15, Corollary 11.17] tells us that the leading term in the asymptotic formulas for both (5.1) and
+we take the sum of (58) over all $\chi$ modulo $q$. Then,
+$\sum_{\chi\,{\rm mod}\,q}f_{\chi}(x)=\varphi(q)\sum_{{n\leq x}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}.$ (72)
+The Dirichlet theorem on arithmetic progressions [15, Corollary 11.17] tells us that the leading term in the asymptotic formulas for both (72) and
 $2\,\sum_{\chi\,{\rm mod}\,q}\sum_{n\leq x}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}=2\,\varphi(q)\sum_{{n\leq x}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}$
 as $x\to\infty$ is $4\sqrt{x}$. Thus, leading terms cancel out in
-$f_{q}(x)$ $=\frac{1}{\varphi(q)}\sum_{\chi\,{\rm mod}\,q}f_{\chi}(x)-\frac{4\sqrt{x}}{\varphi(q)}$ (5.2)
+$f_{q}(x)$ $=\frac{1}{\varphi(q)}\sum_{\chi\,{\rm mod}\,q}f_{\chi}(x)-\frac{4\sqrt{x}}{\varphi(q)}$ (73)
 $=\sum_{{n\leq x}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}-\frac{4\sqrt{x}}{\varphi(q)}$
 and
-$F_{q}(x)$ $=\frac{1}{\varphi(q)}\sum_{\chi\,{\rm mod}\,q}f_{\chi}(x)-\frac{2}{\varphi(q)}\sum_{\chi\,{\rm mod}\,q}\sum_{n\leq x}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}$ (5.3)
+$F_{q}(x)$ $=\frac{1}{\varphi(q)}\sum_{\chi\,{\rm mod}\,q}f_{\chi}(x)-\frac{2}{\varphi(q)}\sum_{\chi\,{\rm mod}\,q}\sum_{n\leq x}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}$ (74)
 $=\frac{1}{\varphi(q)}\sum_{\chi\,{\rm mod}\,q}\sum_{n\leq x}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}\log\frac{x}{ne^{2}}$
 $=\sum_{{n\leq x}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{ne^{2}}.$
-Taking the sum $\varphi(q)^{-1}\sum_{\chi\,{\rm mod}\,q}$ in (4.5),
-$\int_{1}^{\infty}$ $(-f_{q}(x))x^{-s+1/2}\,\frac{dx}{x}=\frac{1}{\varphi(q)}\left(\frac{1}{(s-1/2)^{2}}\sum_{\chi\,{\rm mod}\,q}\frac{L^{\prime}}{L}(s,\chi)+\frac{4}{s-1}\right)$ (5.4)
-for $\Re(s)>1$. On the other hand, by (3.6), (4.5), and the way of their proofs, we obtain
-$\int_{1}^{\infty}\left(-\sum_{n\leq x}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}\log\frac{x}{ne^{2}}\right)x^{-s+1/2}\,\frac{dx}{x}=\frac{2(1-s)}{(s-1/2)^{2}}\frac{L^{\prime}}{L}(s,\chi)$ (5.5)
-for $\Re(s)>1$. Taking the sum $\varphi(q)^{-1}\sum_{\chi\,{\rm mod}\,q}$ in (5.5),
-$\int_{1}^{\infty}(-F_{q}(x))x^{-s+1/2}\,\frac{dx}{x}=\frac{2(1-s)}{(s-1/2)^{2}}\frac{1}{\varphi(q)}\sum_{\chi\,{\rm mod}\,q}\frac{L^{\prime}}{L}(s,\chi).$ (5.6)
-Then, following an argument similar to the proof of Theorems 2 and 8, and using (5.4) and (5.6), we obtain the first claim of Theorem 6.
+Taking the sum $\varphi(q)^{-1}\sum_{\chi\,{\rm mod}\,q}$ in (59),
+$\int_{1}^{\infty}$ $(-f_{q}(x))x^{-s+1/2}\,\frac{dx}{x}=\frac{1}{\varphi(q)}\left(\frac{1}{(s-1/2)^{2}}\sum_{\chi\,{\rm mod}\,q}\frac{L^{\prime}}{L}(s,\chi)+\frac{4}{s-1}\right)$ (75)
+for $\Re(s)>1$. On the other hand, by (51), (59), and the way of their proofs, we obtain
+$\int_{1}^{\infty}\left(-\sum_{n\leq x}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}\log\frac{x}{ne^{2}}\right)x^{-s+1/2}\,\frac{dx}{x}=\frac{2(1-s)}{(s-1/2)^{2}}\frac{L^{\prime}}{L}(s,\chi)$ (76)
+for $\Re(s)>1$. Taking the sum $\varphi(q)^{-1}\sum_{\chi\,{\rm mod}\,q}$ in (76),
+$\int_{1}^{\infty}(-F_{q}(x))x^{-s+1/2}\,\frac{dx}{x}=\frac{2(1-s)}{(s-1/2)^{2}}\frac{1}{\varphi(q)}\sum_{\chi\,{\rm mod}\,q}\frac{L^{\prime}}{L}(s,\chi).$ (77)
+Then, following an argument similar to the proof of Theorems 2 and 8, and using (75) and (77), we obtain the first claim of Theorem 6.
 We now prove the second and third claims of Theorem 6.
 Suppose that $L(1/2,\chi)\not=0$ for any Dirichlet character $\chi$ modulo $q$.
-In an argument similar to the proof of (3.8), in the case $\chi=\chi_{0}$, we include the zeros of $\prod_{p|q}(1-p^{-s})$ in both sums for $\rho$ on the right-hand side of (3.8); in the case $\chi\neq\chi_{0}$, we use (4.8) and (2.13) instead of (3.1) and (2.6), respectively, and include the zeros of $\prod_{p|q}(1-\chi^{\ast}(p)p^{-s})$ in both sums for $\rho$ on the right-hand side, to obtain
-$\sum_{n\leq x}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}\log\frac{x}{ne^{2}}$ $=-\frac{L^{\prime}}{L}\left(\frac{1}{2},\chi\right)\log x+2\sum_{\rho_{\chi}}\frac{x^{\rho_{\chi}-1/2}}{\rho_{\chi}-1/2}$ (5.7)
+In an argument similar to the proof of (53), in the case $\chi=\chi_{0}$, we include the zeros of $\prod_{p|q}(1-p^{-s})$ in both sums for $\rho$ on the right-hand side of (53); in the case $\chi\neq\chi_{0}$, we use (62) and (43) instead of (46) and (36), respectively, and include the zeros of $\prod_{p|q}(1-\chi^{\ast}(p)p^{-s})$ in both sums for $\rho$ on the right-hand side, to obtain
+$\sum_{n\leq x}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}\log\frac{x}{ne^{2}}$ $=-\frac{L^{\prime}}{L}\left(\frac{1}{2},\chi\right)\log x+2\sum_{\rho_{\chi}}\frac{x^{\rho_{\chi}-1/2}}{\rho_{\chi}-1/2}$ (78)
 $\quad-\sum_{\rho_{\chi}}\frac{x^{\rho_{\chi}-1/2}}{(\rho_{\chi}-1/2)^{2}}-\left(\frac{L^{\prime}}{L}\right)^{\prime}\left(\frac{1}{2},\chi\right)+2\frac{L^{\prime}}{L}\left(\frac{1}{2},\chi\right)$
@@
 Here, the two sums $\sum_{\rho_{\chi}}$ on the right-hand side are taken over all nontrivial zeros of $L(s,\chi^{\ast})$ and all zeros of the finite product $L(s,\chi)/L(s,\chi^{\ast})=\prod_{p\mid q}(1-\chi^{\ast}(p)p^{-s})$, $\kappa^{\ast}=\kappa(\chi^{\ast})$, and the last term arises from the difference between $\sideset{}{{}^{\prime}}{\sum}_{n\leq x}$ and $\sum_{n\leq x}$.
-Taking the sum $\sum_{\chi\,{\rm mod}\,q}$ in (4.8) and using (3.1) for $\chi=\chi_{0}$, we obtain
-$\varphi(q)f_{q}(x)$ $=-\log x\sum_{\chi\,{\rm mod}\,q}\frac{L^{\prime}}{L}\left(\frac{1}{2},\chi\right)-\sum_{\chi\,{\rm mod}\,q}\sum_{\rho_{\chi^{\ast}}}\frac{x^{\rho_{\chi^{\ast}}-1/2}}{(\rho_{\chi^{\ast}}-1/2)^{2}}+O(1),$ (5.8)
+Taking the sum $\sum_{\chi\,{\rm mod}\,q}$ in (62) and using (46) for $\chi=\chi_{0}$, we obtain
+$\varphi(q)f_{q}(x)$ $=-\log x\sum_{\chi\,{\rm mod}\,q}\frac{L^{\prime}}{L}\left(\frac{1}{2},\chi\right)-\sum_{\chi\,{\rm mod}\,q}\sum_{\rho_{\chi^{\ast}}}\frac{x^{\rho_{\chi^{\ast}}-1/2}}{(\rho_{\chi^{\ast}}-1/2)^{2}}+O(1),$ (79)
 where the sums $\sum_{\rho_{\chi^{\ast}}}$ on the right-hand side are taken over all nontrivial zeros of $L(s,\chi^{\ast})$ and the implied constant depends only on $q$.
-Similarly, taking the sum $\sum_{\chi\,{\rm mod}\,q}$ in (5.7), we obtain
-$\varphi(q)F_{q}(x)$ $=-\log x\sum_{\chi\,{\rm mod}\,q}\frac{L^{\prime}}{L}\left(\frac{1}{2},\chi\right)$ (5.9)
+Similarly, taking the sum $\sum_{\chi\,{\rm mod}\,q}$ in (78), we obtain
+$\varphi(q)F_{q}(x)$ $=-\log x\sum_{\chi\,{\rm mod}\,q}\frac{L^{\prime}}{L}\left(\frac{1}{2},\chi\right)$ (80)
 $\quad+\sum_{\chi\,{\rm mod}\,q}\left[2\sum_{\rho_{\chi^{\ast}}}\frac{x^{\rho_{\chi^{\ast}}-1/2}}{\rho_{\chi^{\ast}}-1/2}-\sum_{\rho_{\chi^{\ast}}}\frac{x^{\rho_{\chi^{\ast}}-1/2}}{(\rho_{\chi^{\ast}}-1/2)^{2}}\right]+O(1),$
-where the sums $\sum_{\rho_{\chi}^{\ast}}$ and the implied constant have the same meaning as in (5.8).
-In (5.9), the contribution from the zeros of $L(s,\chi)/L(s,\chi^{\ast})$ is included in the $O(1)$ term, and it is justified as follows.
+where the sums $\sum_{\rho_{\chi}^{\ast}}$ and the implied constant have the same meaning as in (79).
+In (80), the contribution from the zeros of $L(s,\chi)/L(s,\chi^{\ast})$ is included in the $O(1)$ term, and it is justified as follows.
 For a Dirichlet character $\chi$ modulo $q$, we denote by $\mathcal{Z}_{1}(\chi)$ the set of all nontrivial zeros of $L(s,\chi^{\ast})$ counted with multiplicity and by $\mathcal{Z}_{2}(\chi)$ the set of all zeros of $L(s,\chi)/L(s,\chi^{\ast})$.
-Then, the second sum on the right-hand side of (2.13) for $s=1/2$ is written as
+Then, the second sum on the right-hand side of (43) for $s=1/2$ is written as
 $\sum_{\rho_{\chi}}\frac{x^{\rho_{\chi}-1/2}}{\rho_{\chi}-1/2}=\lim_{T\to\infty}\left(\sum_{{\rho_{\chi}\in\mathcal{Z}_{1}(\chi)}\atop{|\Im(\rho_{\chi})|\leq T}}\frac{x^{\rho_{\chi}-1/2}}{\rho_{\chi}-1/2}+\sum_{{\rho_{\chi}\in\mathcal{Z}_{2}(\chi)}\atop{|\Im(\rho_{\chi})|\leq T}}\frac{x^{\rho_{\chi}-1/2}}{\rho_{\chi}-1/2}\right).$
@@
 is convergent and bounded, as in the case of the Fourier series of the sawtooth wave.
-Therefore, the contribution from the zeros of $L(s,\chi)/L(s,\chi^{\ast})$ on the right-hand side of (5.7), which originates from (2.13), is bounded.
-Moreover, the contributions from the zeros of $L(s,\chi)/L(s,\chi^{\ast})$ arising from (3.1) and (4.8) are also bounded, since the sums over zeros in (3.1) and (4.8) converge absolutely.
-Assuming that the GRH holds for $L(s,\chi)$ for all $\chi$ modulo $q$, the second sum on the right-hand side of (5.8) is bounded.
-Additionally, if we assume (1.25), the last sum over $\chi$ mod $q$ on the right-hand side of (5.9) is $o(\log x)$.
-Hence, (1.23) and (1.24) hold.
-Conversely, assuming (1.23) (resp.
-(1.24)), it follows from (5.2) (resp.
-(5.3)) that $f_{q}(x)\ll\log x$ (resp.
+Therefore, the contribution from the zeros of $L(s,\chi)/L(s,\chi^{\ast})$ on the right-hand side of (78), which originates from (43), is bounded.
+Moreover, the contributions from the zeros of $L(s,\chi)/L(s,\chi^{\ast})$ arising from (46) and (62) are also bounded, since the sums over zeros in (46) and (62) converge absolutely.
+Assuming that the GRH holds for $L(s,\chi)$ for all $\chi$ modulo $q$, the second sum on the right-hand side of (79) is bounded.
+Additionally, if we assume (25), the last sum over $\chi$ mod $q$ on the right-hand side of (80) is $o(\log x)$.
+Hence, (23) and (24) hold.
+Conversely, assuming (23) (resp.
+(24)), it follows from (73) (resp.
+(74)) that $f_{q}(x)\ll\log x$ (resp.
 $F_{q}(x)\ll\log x$) as $x\to\infty$.
-Therefore, the integral on the left-hand side of (5.4) (resp.
-(5.6)) converges uniformly on any compact subset of the right half-plane $\Re(s)>1/2$, and hence defines an analytic function there.
+Therefore, the integral on the left-hand side of (75) (resp.
+(77)) converges uniformly on any compact subset of the right half-plane $\Re(s)>1/2$, and hence defines an analytic function there.
 Thus, the GRH for $L(s,\chi)$ for all $\chi$ modulo $q$ follows from an argument similar to that in the proof of Theorem 8.
-Moreover, assuming (1.24), it follows from (5.3) that
+Moreover, assuming (24), it follows from (74) that
 $\varphi(q)F_{q}(x)=-\log x\sum_{\chi\,{\rm mod}\,q}\frac{L^{\prime}}{L}\left(\frac{1}{2},\chi\right)+o(\log x)\quad\text{as }x\to\infty.$
-Comparing this with (5.9), we find that the sum over $\chi$ mod $q$ on the right-hand side of (5.9) must be $o(\log x)$.
-However, the latter part of this sum is absolutely convergent and bounded, and thus (1.25) follows.
+Comparing this with (80), we find that the sum over $\chi$ mod $q$ on the right-hand side of (80) must be $o(\log x)$.
+However, the latter part of this sum is absolutely convergent and bounded, and thus (25) follows.
 Finally, we prove the fourth and fifth claims of Theorem 6.
-By using (2.14) in place of (2.13) and (4.10) in place of (4.8), and employing the resulting modified versions of (5.8) and (5.9), one can show that both (1.26) and (1.27) are equivalent to the GRH for $L(s,\chi)$ for all Dirichlet characters $\chi$ modulo $q$, by an argument similar to that used in the proofs of the second and third claims.
-Furthermore, by appropriately modifying (5.9) using (2.14) and (4.10), one can also show that (1.27) implies (1.28) by an argument similar to that of the proof of the third claim.
+By using (44) in place of (43) and (64) in place of (62), and employing the resulting modified versions of (79) and (80), one can show that both (26) and (27) are equivalent to the GRH for $L(s,\chi)$ for all Dirichlet characters $\chi$ modulo $q$, by an argument similar to that used in the proofs of the second and third claims.
+Furthermore, by appropriately modifying (80) using (44) and (64), one can also show that (27) implies (28) by an argument similar to that of the proof of the third claim.
 We omit the details, as there is no essential difference.
@@
 5.2. Proof of Theorem 7
-First, we prove (1.29). After multiplying $f_{q}(x)$ in (5.2) by $\varphi(q)$, summing over $3\leq q\leq Q$, and then changing the order of summation, we obtain
-$\sum_{3\leq q\leq Q}\varphi(q)f_{q}(x)$ $=\sum_{3\leq q\leq Q}\left(\varphi(q)\sum_{{n\leq x}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}-4\sqrt{x}\right)$ (5.10)
+First, we prove (29). After multiplying $f_{q}(x)$ in (73) by $\varphi(q)$, summing over $3\leq q\leq Q$, and then changing the order of summation, we obtain
+$\sum_{3\leq q\leq Q}\varphi(q)f_{q}(x)$ $=\sum_{3\leq q\leq Q}\left(\varphi(q)\sum_{{n\leq x}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}-4\sqrt{x}\right)$ (81)
 $=\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}\sum_{{3\leq q\leq Q}\atop{q\mid n-1}}\varphi(q)-4\sqrt{x}\,(Q-2),$
@@
 For a prime number $p$, 22 divides $p^{k}-1$ if and only if $p$ is odd.
-Therefore, using $\sum_{d|m}\varphi(d)=m$ on the right-hand side of (5.10),
+Therefore, using $\sum_{d|m}\varphi(d)=m$ on the right-hand side of (81),
 $\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}\sum_{{3\leq q\leq Q}\atop{q\mid n-1}}\varphi(q)$ $=\sum_{{n\leq x}\atop{n\not\in 2^{\mathbb{N}}}}\frac{\Lambda(n)(n-3)}{\sqrt{n}}\log\frac{x}{n}+\log 2\sum_{2^{k}\leq x}\frac{(2^{k}-2)}{\sqrt{2^{k}}}\log\frac{x}{2^{k}}$
 $=\sum_{n\leq x}\frac{\Lambda(n)(n-3)}{\sqrt{n}}\log\frac{x}{n}+o(\sqrt{x}\,).$
-Using [21, (2.3)] and the prime number theorem (2.2) in (3.1),
+Using [21, (2.3)] and the prime number theorem (32) in (46),
 $\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}=O(\sqrt{x}).$
@@
 $\sum_{3\leq q\leq Q}\varphi(q)f_{q}(x)=\sum_{n\leq x}\Lambda(n)\sqrt{n}\,\log\frac{x}{n}-4\sqrt{x}\,Q+O(\sqrt{x}\,).$
-By applying partial summation to the first term on the right-hand side using the prime number theorem (2.2),
+By applying partial summation to the first term on the right-hand side using the prime number theorem (32),
 $\sum_{n\leq x}\Lambda(n)\cdot\sqrt{n}\,\log\frac{x}{n}$ $=-\int_{1}^{x}\sum_{n\leq u}\Lambda(n)\left(-\frac{1}{\sqrt{u}}+\frac{1}{2\sqrt{u}}\log\frac{x}{u}\right)\,du$
 $=\frac{4}{9}x\sqrt{x}\,(1+o(1)).$
-Hence, we obtain (1.29).
-Next, we prove (1.30). After multiplying $F_{q}(x)$ in (5.3) by $\varphi(q)$, summing over $3\leq q\leq Q$, and then changing the order of summation as in (5.10), we obtain
-$\sum_{3\leq q\leq Q}\varphi(q)F_{q}(x)=\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{ne^{2}}\sum_{{3\leq q\leq Q}\atop{q\mid n-1}}\varphi(q).$ (5.11)
+Hence, we obtain (29).
+Next, we prove (30). After multiplying $F_{q}(x)$ in (74) by $\varphi(q)$, summing over $3\leq q\leq Q$, and then changing the order of summation as in (81), we obtain
+$\sum_{3\leq q\leq Q}\varphi(q)F_{q}(x)=\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{ne^{2}}\sum_{{3\leq q\leq Q}\atop{q\mid n-1}}\varphi(q).$ (82)
 Then, as before,
 $\sum_{3\leq q\leq Q}\varphi(q)F_{q}(x)=\sum_{n\leq x}\frac{\Lambda(n)(n-3)}{\sqrt{n}}\log\frac{x}{ne^{2}}+o(\sqrt{x}\,).$
-Using [21, (2.3)] and the prime number theorem (2.2) in (3.8),
+Using [21, (2.3)] and the prime number theorem (32) in (53),
 $\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{ne^{2}}=o(\sqrt{x}\,).$
@@
 $\sum_{3\leq q\leq Q}\varphi(q)F_{q}(x)=\sum_{n\leq x}\Lambda(n)\sqrt{n}\,\log\frac{x}{ne^{2}}+o(\sqrt{x}\,).$
-By applying partial summation to the first term on the right-hand side using the prime number theorem (2.2),
+By applying partial summation to the first term on the right-hand side using the prime number theorem (32),
 $\sum_{n\leq x}\Lambda(n)\cdot\sqrt{n}\,\log\frac{x}{ne^{2}}$ $=-2\sqrt{x}\,\sum_{n\leq x}\Lambda(n)-\int_{1}^{x}\sum_{n\leq u}\Lambda(n)\left(-\frac{2}{\sqrt{u}}+\frac{1}{2\sqrt{u}}\log\frac{x}{u}\right)\,du$
 $=-\frac{8}{9}x\sqrt{x}\,(1+o(1)).$
-Hence, we obtain (1.30) by substituting $xe^{2}$ for $x$. $\Box$
-If we take the sum $\sum_{3\leq q\leq Q}$ in (5.3), we obtain
+Hence, we obtain (30) by substituting $xe^{2}$ for $x$. $\Box$
+If we take the sum $\sum_{3\leq q\leq Q}$ in (74), we obtain
 $\sum_{3\leq q\leq Q}F_{q}(x)=\sum_{n\leq x}\frac{\Lambda(n)(\sigma_{0}(n-1)-\nu(n))}{\sqrt{n}}\log\frac{x}{ne^{2}},$
@@
 $\sum_{n\leq x}\frac{\Lambda(n)\sigma_{0}(n-1)}{\sqrt{n}}\log\frac{x}{ne^{2}}=-8\,\frac{\zeta(2)\zeta(3)}{\zeta(6)}\sqrt{x}\,(1+o(1)).$
-This also shows that (1.22) has a constant sign on average, however, the sum used in the proof of Theorem 7 is probably easier to handle.
+This also shows that (22) has a constant sign on average, however, the sum used in the proof of Theorem 7 is probably easier to handle.